### v2

# From Restless to Contextual: A Thresholding Bandit Approach to Improve Finite-horizon Performance

Abstract
Online restless bandits extend classic contextual bandits by incorporating state transitions and budget constraints, representing each agent as a Markov Decision Process (MDP). This framework is crucial for finite-horizon strategic resource allocation, optimizing limited costly interventions for long-term benefits. However, learning the underlying MDP for each agent poses a major challenge in finite-horizon settings. To facilitate learning, we reformulate the problem as a scalable budgeted thresholding contextual bandit problem. We carefully integrate the state transitions into the reward design and focus on identifying agents with action benefits exceeding a threshold. We establish the optimality of an oracle greedy solution in a simple two-state setting, and propose an algorithm that achieves minimax optimal constant regret in the online multi-state setting with heterogeneous agents and knowledge of outcomes under no intervention. We numerically show that our algorithm outperforms existing online restless bandit methods, offering significant improvements in finite-horizon performance.
Online Restless Bandit, Contextual Threshold Bandit, Regret Minimization, Finite-horizon Optimization
1 Introduction
Restless bandits extend classic contextual bandits by modeling state transitions and incorporating budget constraints, providing a framework for dynamic decision-making in nonstationary environments. In this framework, a set of independent arms (agents) evolves over discrete time steps according to a Markov Decision Process (MDP), regardless of whether they are actively played or passively rested. The goal is to dynamically allocate a fixed number of active plays to maximize cumulative rewards over time. While much of the existing literature focuses on offline settings, where transitions and rewards are assumed known (Whittle, 1988; Hong et al., 2024), online restless bandits, where these parameters must be learned, have gained traction due to their practical relevance (Jung & Tewari, 2019).
This framework is well-suited for real-world applications where interventions are costly, resources are limited, and outcomes evolve over time. For example, in healthcare, restless bandits have been used to optimize targeted interventions (Villar et al., 2015; Biswas et al., 2021), such as guidance calls in mobile health programs to improve engagement in automated prenatal care (Mate et al., 2021). They have also been applied to resource allocation problems (Borkar et al., 2017) and other adaptive decision-making (Brown & Smith, 2020). In practice, interventions are often rolled out gradually, or there is abundant observational data from periods without intervention (Liu et al., 2014; Shaikh et al., 2019; Mate et al., 2021; Periáñez et al., 2024), enabling the estimation of state evolution dynamics in the absence of intervention. Further, these settings often operate under finite horizons, where the experimentation period for new interventions is constrained by factors such as budget limitations, client-imposed timeframes, evolving market conditions, or changing health guidelines (del Río et al., 2024a, b). Table 1 provides a broader overview of real-world applications where restless bandits naturally arise.
Table 1: Examples of good and bad states and corresponding budgeted interventions in various domains
Domain Good state Bad state Budgeted Intervention Example
General Usage Frequent logins, feature exploration, meaningful use Rare/brief logins, session abandonment Reach out, offer help, gather feedback
E-commerce High purchase frequency, browsing multiple categories Low/no purchases, high cart abandonment Provide a personalized discount or free shipping
E-learning Completes modules, passes quizzes, revisits content Seldom opens modules, fails quizzes Progress‐based reward (finish by deadline for free certificate)
Loyalty Completes surveys, redeems rewards, advances loyalty Not collecting points, ignoring surveys, no rewards Special bonus for feedback (earn points by reviewing new features)
Patient management Updates records, completes follow‐ups, tracks adherence Rarely logs updates, misses assessments or follow‐ups Points or bonus for timely follow‐ups
Despite their potential, solving online restless bandits in finite-horizon settings is challenging, as they require detailed knowledge of the underlying MDP for each agent (Jung & Tewari, 2019; Wang et al., 2020, 2023b). In such settings, this leads to an unavoidable linear regret when maximizing cumulative rewards (Azar et al., 2017). These difficulties call for a reformulation of the problem to simplify the learning objective, enabling the development of solutions that are not only optimal for the reformulated problem but also demonstrate improved empirical performance in finite samples, even in the online restless bandit setting.
Our contributions are three-fold. First, we reformulate online restless bandits as a budgeted thresholding contextual bandit, simplifying learning through carefully designed rewards. This reformulation focuses on identifying agents whose action benefits exceed a predefined threshold at each time step (Section 2). We establish the optimality of this reformulation in a simple 2-state, finite horizon setting (Theorem 3.6) where agents are homogeneous and giving intervention is always better than no intervention (Assumption 3.1). We note that this simple setting has been applied to many real-world settings (Wang et al., 2023a; Killian et al., 2023; Liang et al., 2024; Raman et al., 2024), highlighting the utility of our reformulation.
Second, assuming knowledge of outcomes under no intervention, we propose a novel learning algorithm for heterogeneous agents in multi-state settings: the lower confidence bound (LCB)-guided randomized thresholding algorithm (LGRT, Algorithm 1, Section 4). We establish that LGRT achieves optimal minimax constant regret under our reformulation (Theorem 4.4), enabled by a novel design of the LCB for the reward empirical mean.
Lastly, we numerically demonstrate (in Section 5) that a greedy variant of our proposed algorithm, Algorithm 2, consistently outperforms existing online restless bandit algorithms in finite-horizon online restless bandit settings for instances satisfying stochastic dominance assumption (Assumption 3.1), validating its effectiveness and practical utility. We further evaluate the performance of our algorithm in the absence of the stochastic dominance assumption. We modify both our algorithm and the benchmarks, allowing them to allocate actions to fewer than B agents when additional actions are expected to yield negative benefits. Our numerical results demonstrate that our modified algorithm, Algorithm 3, outperforms the adapted benchmarks, highlighting the practicality of our framework.
1.1 Related Work
The offline restless bandit problem, with known transition matrices and reward functions, is PSPACE-hard to solve due to the combinatorial nature of the budget constraint (Papadimitriou & Tsitsiklis, 1999). Many approximation algorithms have been proposed, including Whittle’s index policy (Whittle, 1988; Weber & Weiss, 1990; Hsu, 2018; Kadota et al., 2016) and linear programming-based methods (Verloop, 2016; Zhang & Frazier, 2021; Brown & Zhang, 2023; Hong et al., 2024; Gast et al., 2024), both of which are proven to be optimal as the number of agents approaches infinity. Whittle’s index policy is commonly used as a benchmark for evaluating online restless bandit algorithms (Jung & Tewari, 2019; Wang et al., 2020, 2023b; Akbarzadeh & Mahajan, 2023). We provide a brief overview in Section 2.1 and a detailed explanation in Appendix A. In this paper, we focus on the online setting, where transitions and rewards must be learned from data.
Online Restless Bandit   Online restless bandit algorithms typically fall into two categories: model-based and model-free methods. The former directly estimates the transition matrix in both finite-(Jung & Tewari, 2019; Wang et al., 2023b; Liang et al., 2024) and infinite-horizon (Ortner et al., 2012; Wang et al., 2020; Xiong et al., 2022b, a; Akbarzadeh & Mahajan, 2023) settings, whereas the latter focuses on learning the Q-function specifically in the infinite-horizon setting (Killian et al., 2021; Xiong & Li, 2023; Xiong et al., 2024; Kakarapalli et al., 2024). Our problem focuses on the finite horizon setting, and we contrast our results with prior finite-horizon regret guarantees in Section 2.2.
Thresholding Bandit/MDP   Our paper is closely related to regret-minimizing thresholding bandit problems, where regret is defined by the gap between the arm’s mean and a predefined threshold when the mean falls below the threshold. It is shown that under this setting, minimax optimal (max is taken over all possible bandit instances and min taken over all possible learning algorithms) constant regret is achievable (Tamatsukuri & Takahashi, 2019; Michel et al., 2022; Feng et al., 2024). Our problem setup differs from the traditional thresholding bandit framework by incorporating context, where the arrival of contexts follows an MDP. In contrast, Hajiabolhassan & Ortner (2023) study threshold MDPs under an infinite horizon setting, seeking policies with average rewards exceeding a predefined threshold. We, however, focus on the finite-horizon setting.
Episodic MDP   When both the budget and the number of arms in an online restless bandit are set to one, the problem reduces to a standard reinforcement learning problem. Thus, our work also relates to episodic MDP, which focuses on maximizing cumulative rewards across a finite horizon. Let $K$ be the number of episodes, each containing $T$ time steps. Azar et al. (2017) establishes a minimax regret of $\mathcal{O}\left(\sqrt{T^{2}K}\right)$ for stationary transitions, and Jin et al. (2018) derives an upper bound of $\mathcal{O}\left(\sqrt{T^{4}K}\right)$ for nonstationary transitions. These results highlight the difficulty of the learning problem under a finite horizon, emphasizing the need for a reformulation to simplify the learning objective.
2 Restless Bandits and the Budgeted Thresholding Contextual Bandit Reformulation
This section begins by formally defining the restless bandit problem with known transition probabilities and reward function in Section 2.1. In Section 2.2, we extend the problem to the online setting, where transition and reward distributions are unknown and must be learned from data. We note that online restless bandit learning is data-intensive, making it impractical for budget-constrained settings with a finite horizon. To address this limitation, Section 2.3 presents our problem reformulation, and Section 3 characterizes the conditions under which this reformulation achieves optimality for the restless bandit problem in the offline setting.
2.1 Restless Bandits
In an offline finite-horizon restless bandit problem, the learner is provided with $M$ independent arms, each representing one independent agent. At each time $t\in[T]=\{1,\ldots,T\}$, each arm, $m\in[M]$, is associated with a state $s_{t}^{m}\in\mathcal{S}$, and an action $a_{t+1}^{m}\in\mathcal{A}$. We focus on binary action, i.e., $\mathcal{A}=\{0,1\}$. Each agent $m$ is modeled as an MDP represented by the tuple $\left(\mathcal{S},\mathcal{A},P_{a}^{m},R^{m}\right)$, where $P_{a}^{m}(s,s^{\prime}):\mathcal{S}\times\mathcal{S}\to[0,1]$ represents the transition probability from state $s$ to $s^{\prime}$ under action $a$ for agent $m$, and $R^{m}(s):\mathcal{S}\to\mathbb{R}$ represents the expected reward that agent $m$ receives at state $s$. We highlight that in a restless bandit problem, the expected reward, $R^{m}(s)$, is independent of the action, and the learner observes the reward immediately after arriving at the state. In the offline setting, both $P_{a}^{m}(s,s^{\prime})$ and $R^{m}(s)$ are known to the learner.
At each time $t$, the learner observes the set of states of $M$ agents, denoted by $\bm{s}_{t}:=\left(s_{t}^{1},s_{t}^{2},\cdots,s_{t}^{M}\right)$, and receives the corresponding reward $\bm{r}_{t}:=(r_{t}^{1},\cdots,r_{t}^{M})$. Based on the observed states, the learner determines the action $\bm{a}_{t+1}:=\left(a_{t+1}^{1},a_{t+1}^{2},\cdots,a_{t+1}^{M}\right)$, according to a deterministic policy $\pi^{d}_{t+1}(\bm{s}):\mathcal{S}^{M}\rightarrow\{0,1\}^{M}$. The learner is provided with a budget constraint, allowing at most $B$ arms to be pulled at each time step, i.e., $\sum_{m\in[M]}a_{t+1}^{m}\leq B$. After pulling up to $B$ arms, the learner observes the next state $\bm{s}_{t+1}$, generated according to $s_{t+1}^{m}\sim P_{a_{t+1}^{m}}^{m}\left(s_{t}^{m},s_{t+1}^{m}\right)$, and receives a feedback $\bm{r}_{t+1}$.
Let $\bm{s}_{0}$ denote the initial states and $\bm{r}_{0}$ their associated rewards. The learner aims to find the policy $\bm{\pi}_{d}:=\{\pi^{d}_{t}\}_{t\in[T]}$ that maximizes the expected cumulative reward,
$\max_{{\bm{\pi}}_{d}}\;$ $\mathbb{E}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]$
s.t. $\sum_{m=1}^{M}a_{t}^{m}\leq B,\;\forall t\in[T],$ (1)
where the expectation is taken over ${s}_{t+1}^{m}\sim P_{a}^{m}(s_{t}^{m},s_{t+1}^{m})$ for all $m$. We will use $\bm{\pi}^{*}$ to denote the optimal policy corresponding to Problem (1).
Whittle’s Index Policy   This policy provides an approximate solution to Problem (1) by maximizing the average expected reward as the horizon length approaches infinity (A.9). [^1] Alternatively, Whittle’s index policy can also be defined for maximizing the discounted reward as the horizon approaches infinity (A.11), which we describe in Appendix A. It incorporates the budget constraint using Lagrangian relaxation (A.10), allowing the problem to be decoupled and solved individually for each agent. Whittle’s index policy achieves the optimal solution to Problem (A.9) as the number of homogeneous agents $M$ approaches infinity (Theorem A.1). Let $\bm{\pi}^{\text{Whittle}}$ be Whittle’s index policy obtained in the offline setting, corresponding to the solution to Eq. (A.10). In Section 5, we benchmark the performance of our algorithm against that of $\bm{\pi}^{\text{Whittle}}$ in our problem instances. We describe Whittle’s index policy in detail in Appendix A.
2.2 Online Restless Bandits
In this setting, the transition matrices (and/or reward functions) need to be learned from the data. As a result, at each time $t$, given the state $\bm{s}_{t}$, the learner’s decision $\bm{a}_{t+1}$ follows a history-dependent random policy, $\pi_{t+1}$. Let $\mathcal{H}_{t}:=\{(\bm{a}_{i},\bm{s}_{i},\bm{r}_{i})\}_{i=1}^{t}\cup\{\bm{s}_{0},\bm{r}_{0}\}$ be the history observed up to time $t$, and let $\mathcal{H}_{t}^{+}:=\mathcal{H}_{t}\cup\{\bm{a}_{t+1}\}$ be the augmented history up to taking action $\bm{a}_{t+1}$. Then, $\pi_{t+1}:\mathcal{H}_{t}\to\Delta{\left(\{0,1\}^{M}\right)}$, where $\Delta{\left(\{0,1\}^{M}\right)}$ denotes the probability simplex over all possible actions. Suppose that the reward $\bm{r}_{t+1}$ is generated according to a distribution $\mathcal{P}_{t}$, we will assume stationarity of the expected reward received at time $t+1$, $\bm{r}_{t+1}$, when conditioned on history $\mathcal{H}_{t}^{+}$ and the current state.
Assumption 2.1.
$\mathbb{E}\left[r_{t+1}^{m}\mid s_{t+1}^{m},\mathcal{H}_{t}^{+}\right]=R^{m}\left(s_{t+1}^{m}\right).$
We permit non-stationarity in the rewards, provided that Assumption 2.1 is satisfied, offering a relaxation of the standard stationarity assumption typically imposed in online restless bandit problems.
Similar to the offline setting, the goal of the learner is to find the policy $\bm{\pi}:=\{\pi_{t}\}_{t\in[T]}$ that maximizes the expected cumulative reward, where the expectation is further taken over the randomized policy ($\bm{a}_{t}\sim\pi_{t}$), in addition to the states ($s_{t+1}^{m}\sim P^{m}_{a_{t+1}^{m}}(s_{t}^{m},s_{t+1}^{m})$):
$\vspace{-5pt}\max_{{\bm{\pi}}}\;$ $\mathbb{E}_{\bm{\pi}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]$
s.t. $\sum_{m=1}^{M}a_{t}^{m}\leq B,\;\forall t\in[T].$ (2)
Online Whittle’s Index Policy and Hardness of the Finite-Horizon Problem   The online Whittle’s index policy often involves estimating the uncertainty associated with the transition probabilities from data and using it to obtain an online Whittle’s index policy (Akbarzadeh & Mahajan, 2023; Wang et al., 2023b). However, in a finite horizon setting, the challenge of learning transition probabilities becomes even more complex due to the budget constraints. To formally highlight the difficulty of the learning problem within the online restless bandit framework, we examine the regret of the online Whittle’s index policy as an illustrating example. Given a horizon length of $T$ and an online Whittle’s index policy $\bm{\omega}$, the regret $\mathcal{R}_{f}(\bm{\omega})$ is defined as follows: [^2] Oftentimes, in online restless bandits, the horizon is segmented into episodes, each with a fixed horizon length $T$. A generalized definition of regret, along with its discussion, is included in Appendix A.1.
$\mathcal{R}_{f}(\bm{\omega}):=\mathbb{E}_{\bm{\pi}^{\text{Whittle}}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]-\mathbb{E}_{\bm{\omega}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right].$
Wang et al. (2023b) establishes an upper bound of $\mathcal{O}(T^{3/2})$ on $\mathcal{R}_{f}(\bm{\omega})$, and Azar et al. (2017) establishes a lower bound of $\Omega(T)$ when $B=M=1$. This highlights that a sublinear regret in the online restless bandit problem is impossible to achieve, highlighting the difficulty of this learning problem.
2.3 Budgeted Thresholding Contextual Bandit Reformulation
Thus motivated, we propose to simplify the learning objective by making the following modifications:
  1. a)
    We reformulate the problem as a contextual bandit problem, where the transition probabilities no longer need to be estimated explicitly;
  2. b)
    Instead of finding the optimal policy that maximizes the expected cumulative rewards across $M$ agents over a horizon $T$, we focus on identifying arms (agents) at each time step where taking action (versus not) yields an effect size exceeding a predefined threshold.
We observe that a) (to be introduced below) simplifies the learning problem by embedding the nonstationarity of the state in online restless bandits into the reward structure. Consequently, the resulting policy depends solely on the current context. In general, learning a contextual bandit requires significantly lower sample complexity compared to an MDP (Simchi-Levi & Xu, 2022). Additionally, Modification b) further simplifies the problem by reducing reward maximization to a threshold identification task, which is known to have a lower sample complexity (Feng et al., 2024). We establish that under suitable assumptions, our modification, under the knowledge of the transition and reward, recovers the optimal policy of the offline restless bandit problem in Section 3 for any $M$. In contrast, Whittle’s index policy achieves optimality only when the number of agents and horizon length tend to infinity in the offline setting.
Regret-Minimizing Budgeted Thresholding Contextual Bandit Reduction   In the reduced contextual bandit problem, the context space is the state space. At time $t$, the reward $\phi^{m}$ that agent $m$ receives at state $s_{t}^{m}$ under action $a_{t+1}^{m}$ is defined as the expected reward that the agent will receive in the next time period:
$\phi^{m}(s_{t}^{m},a_{t+1}^{m}):=\sum_{s_{t+1}^{m}\in\mathcal{S}}P^{m}_{a_{t+1}^{m}}(s_{t}^{m},s_{t+1}^{m})R^{m}\left(s_{t+1}^{m}\right)$
$\quad\quad=\sum_{s_{t+1}^{m}\in\mathcal{S}}\mathbb{P}\left(s_{t+1}^{m}\mid\mathcal{H}_{t}^{+}\right)\mathbb{E}\left[r_{t+1}^{m}\mid s_{t+1}^{m},\mathcal{H}_{t}^{+}\right]$ (3)
$\quad\quad=\mathbb{E}\left[r_{t+1}^{m}\mid\mathcal{H}_{t}^{+}\right].$ (4)
Eq. (3) follows from Assumption 2.1 and the Markovian property of the transition probabilities. Eq. (4) shows that the immediate reward, $\phi^{m}(s_{t}^{m},a_{t+1}^{m})$, can be estimated without explicitly modeling the transition probabilities. It does not depend on the next state, consistent with the standard bandit reward formulation.
As in the online restless bandits, at each time $t$, the learner observes $\bm{s}_{t}$ and the corresponding reward $\bm{r}_{t}$. Using these observations, the learner updates the estimated uncertainty associated with the estimated $\phi^{m}(s_{t-1}^{m},a_{t}^{m})\;\forall m\in[M]$ according to Eq. (4), and uses it to guide the action $\bm{a}_{t+1}$, with $\sum_{m\in[M]}a_{t+1}^{m}\leq B$.
The goal of the learner is to learn a policy that identifies and selects $B$ agents whose expected gain from taking the action (i.e., $a_{t+1}^{m}=1$) exceeds a predefined threshold in their current state $s_{t}^{m}$. To quantify this expected gain in state $s\in\mathcal{S}$, we define $I^{m}(s):=\phi^{m}(s,1)-\phi^{m}(s,0)$, representing the incremental reward an agent receives by taking the action versus not at state $s$. Let $\gamma>0$ be a predefined threshold. At time $t$, we say that an agent $m$ is in a good state if $I^{m}(s_{t}^{m})\geq\gamma$.
At each time $t$, let $G_{t}:=\{m\in[M]:I^{m}(s_{t}^{m})\geq\gamma\}$ denote the set of agents currently in a good state. In this budgeted thresholding contextual bandit formulation, the optimal policy, $\bm{\pi}^{b}$, involves randomly selecting $B$ agents from $G_{t}$ for intervention when $|G_{t}|\geq B$ and selecting $B$ agents according to the ordering of incremental reward otherwise. For any policy $\bm{\pi}$, potentially learned through an online algorithm, the regret $\mathcal{R}(\bm{\pi})$ is defined as:
$\mathcal{R}(\bm{\pi}):=\mathbb{E}_{\bm{\pi}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}c_{t}^{m}\left(a_{t+1}^{m}\right)|\bm{s}_{0}\right]-\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}c_{t}^{m}\left(a_{t+1}^{m}\right)|\bm{s}_{0}\right],$ (5)
where the expectation is taken with respect to $\bm{a}_{t+1}\sim\pi_{t},s_{t+1}^{m}\sim P_{a_{t+1}^{m}}^{m}(s_{t}^{m},s_{t+1}^{m})$, and
$c_{t}^{m}\left(a_{t+1}^{m}\right):=\left(\gamma-I^{m}(s_{t}^{m})\right)\cdot\mathbbm{1}\left\{I^{m}(s_{t}^{m})<\gamma\right\}\cdot\mathbbm{1}\left\{a_{t+1}^{m}=1\right\}.$ (6)
Eq. (6) shows that a loss is incurred at each step only when an action is taken on an agent who is not in a good state. The incurred cost is determined by the difference between the threshold and the incremental reward, given by $\gamma-I^{m}(s_{t}^{m})$. The goal of the learner is to minimize the cumulative regret, defined in Eq. (5), within a fixed horizon $T$.
We note that while learning algorithms can achieve sublinear regret in the reformulated thresholding contextual bandit setting, this does not guarantee an improved regret bound for the online restless bandit problem, as the $\bm{\pi}^{b}$ may be suboptimal. This highlights a fundamental trade-off between learnability and optimality, further explored in Remark 4.6, and empirically examined in Section 5.
Here, regret serves as a measure of the speed of convergence for the designed policy. As established in Subsection 2.2, the online finite-horizon restless bandit problem inherently suffers from linear regret, meaning any algorithm aiming for optimality (in the homogeneous agent setting and asymptotically) will necessarily exhibit slow convergence. To achieve better finite-sample performance, i.e., faster convergence to a policy, we must relax the requirement for optimality, accepting that the policy might be suboptimal. In this context, we hypothesize that an algorithm designed for maximizing reward in the simplified contextual bandit setting could outperform slower-converging algorithms that target near-optimal performance. This is particularly relevant since the optimality of the online Whittle index policy remains unestablished in the heterogeneous agent setting. We numerically validate this hypothesis in Section 5 and confirm that it holds true.
Before establishing the sublinear regret of our proposed algorithm for the thresholding contextual bandit problem in the heterogeneous agent case (Section 4), we first demonstrate that this thresholding bandit formulation is optimal in a 2-state setting with homogeneous agents under the assumption of stochastic dominance in the next section.
3 Optimality of Budgeted Thresholding Contextual Bandit Reduction
In this section, we assume binary states and we further assume that all agents share a good state with higher $R^{m}$. Without loss of generality (WLOG), we will use 1 to denote the good state and 0 for the bad state. We establish that given oracle knowledge of the transition probabilities and rewards, a natural greedy policy is optimal for the offline restless bandit problem described in Section 2.1 under two additional assumptions: 1) stochastic dominance (Assumption 3.1) and 2) homogeneity of the agents. We note that the optimality of this natural greedy policy indicates that the contextual bandit provides an exact reformulation to the offline restless bandit problem under these conditions.
We first describe the greedy policy. WLOG, suppose that at each time step $t$, the incremental rewards are ordered as $I^{1}\left(s_{t}^{1}\right)\geq I^{2}\left(s_{t}^{2}\right)\geq\cdots\geq I^{M}\left(s_{t}^{M}\right)$. The corresponding greedy policy $\pi_{t}^{\text{greedy}}:\mathcal{S}\rightarrow\{0,1\}^{M}$ is to give actions to the first $B$ agents. Formally, with a slight abuse of notation, let $\pi_{t}^{\text{greedy}}({s}_{t}^{i})$ be the $i$th entry of $\pi_{t}^{\text{greedy}}$, defined as:
$\pi_{t}^{\text{greedy}}({s}_{t}^{i})=\begin{cases}1&i\leq B\\ 0&i>B\end{cases}.$
Let $\bm{\pi}^{\text{greedy}}:=\{\pi_{t}^{\text{greedy}}\}_{t\in[T]}$. We note that $\bm{\pi}^{\text{greedy}}=\bm{\pi}^{b}$ with suitable choices of $\gamma$. A trivial choice of $\gamma$ would be $\max_{m,s}I^{m}(s)$. Next, to establish the optimality of $\bm{\pi}^{\text{greedy}}$ in the offline restless bandit setting, we introduce the stochastic dominance assumption below:
Assumption 3.1 (Stochastic Dominance).
Assume that for any non-decreasing function $f(s)$ of $s\in\mathcal{S}$, the following holds for any $s\in\mathcal{S},m\in[M]$:
$\mathbb{E}_{s^{\prime}\sim P_{1}^{m}(s,s^{\prime})}\left[f(s^{\prime})\right]\geq\mathbb{E}_{s^{\prime}\sim P_{0}^{m}(s,s^{\prime})}\left[f(s^{\prime})\right].$ (7)
Before interpreting Assumption 3.1, we provide a necessary and sufficient condition for its validity in the two-state case.
Lemma 3.2.
Assumption 3.1 holds if and only if $P^{m}_{1}(s,1)\geq P^{m}_{0}(s,1)$ for any $s\in\mathcal{S},m\in[M]$.
Lemma 3.2 (proof in Appendix C.1) indicates that when there are two states, Assumption 3.1 is equivalent to the condition that the transition probability to state 1 under action 1 is always higher (better) than under action 0.
Intuitively, Assumption 3.1 imposes a structure on the transitions to ensure that giving an intervention is consistently more beneficial than not giving one in the future. This implies that in a system with a single agent and two states, the optimal solution is to always provide intervention at every time step, which aligns with $\bm{\pi}^{\text{greedy}}$. This observation is formalized in Lemma 3.3 (proof in Appendix C.2).
Lemma 3.3.
Assuming $B=M=1$, and under Assumption 3.1, $\bm{\pi}^{\text{greedy}}=\bm{\pi}^{*}$.
Next, we establish the optimality of the greedy policy in the case of multiple agents. First, in Lemma 3.4 (proof in Appendix C.3), we show that under Assumption 3.1, utilizing the entire budget is always beneficial.
Lemma 3.4.
Given a budget $B\in[M]$, and assuming Assumption 3.1, the optimal policy $\bm{\pi}^{*}$ selects exactly $B$ agents to take action $1$, ensuring full utilization of the budget.
Remark 3.5 (The Stochastic Dominance Assumption).
Assumption 3.1 holds across a wide range of real-world applications, some of which are illustrated in Table 1. Meanwhile, it was also adopted by Wang et al. (2023b) in their experimental section, which we thoroughly compare against in Section 5. Further, in the offline restless bandit literature, the budget constraint often directly requires the learner to spend all the budget (Whittle, 1988), which aligns with Assumption 3.1. When Assumption 3.1 is violated—such as when $P_{0}^{m}(s,s^{\prime})$ stochastically dominates $P_{1}^{m}(s,s^{\prime})$—the optimal policy may favor not intervening at all.
Building on Lemmas 3.2-3.4, Theorem 3.6 (proof in Appendix C.4) establishes the optimality of the greedy policy under homogeneous agents.
Theorem 3.6.
Assuming all agents are homogeneous, under Assumption 3.1, $\bm{\pi}^{\text{greedy}}=\bm{\pi}^{*}$ for any $0\leq B\leq M<\infty$.
Theorem 3.6 generally does not extend to heterogeneous agents, as demonstrated by the counterexample in Appendix B. Nevertheless, we conjecture that under some additional assumptions, Theorem 3.6 could be approximately extended to the multi-state case, where the greedy policy may achieve a constant-factor approximation of the optimal policy.
Remark 3.7 (Significance of Theorem 3.6).
To our knowledge, this is the first result that establishes the optimality of a policy that holds for any number of agents in an offline finite-horizon restless bandit. Though we assume that all agents are homogeneous, this is common in offline restless bandit problems (Hong et al., 2024; Whittle, 1988; Zhang & Frazier, 2021). Meanwhile, in real-world scenarios, such as with the ARMMAN dataset, researchers often simplify the problem by treating it as a two-state system (Liang et al., 2024; Killian et al., 2023; Wang et al., 2023a).
Before we end this section, we outline a proof sketch for Theorem 3.6. The proof consists of two steps: 1) We reformulate the problem as an MDP where the state space is the Cartesian product of the individual state space, i.e., $\mathcal{S}^{M}$; the action space is the subset of $\{0,1\}^{M}$ that satisfies the budget constraint, and the transition matrix is the product of $P^{m}$ as all agents are independent. 2) We show that the proposed greedy policy $\pi_{t}^{\text{greedy}}$ satisfies the Bellman equation (Eq. (C.35)) for all $t\leq T$ by using backward induction. We prove the induction hypothesis by carefully controlling the difference of the value function (Eq. (C.15)).
4 Budgeted Thresholding Contextual Bandit
In this section, we will present a learning algorithm for the budgeted thresholding contextual bandit reformulation described in Section 2.3, Algorithm 1, in a general setting. We establish that under this reformulation, our algorithm achieves the optimal minimax constant regret. To ease the analysis, in this section, we assume that at each time $t$, there are sufficient agents whose ground truth incremental reward is above the threshold (Assumption 4.2). This assumption is merely used to simplify the proof. In practice, Algorithm 1 can be executed with any chosen value of $\gamma$ and serve as a tunable parameter that can be used to improve the overall algorithm performance. We will discuss this in detail in Remark 4.6. In Section 5, we do not require $\gamma$ to satisfy Assumption 4.2. We assume access to an accurate estimator of the reward $\phi^{m}$ under no intervention, based on the observation that the system’s dynamics are more easily observed without intervention. This does not require directly estimating transition probabilities.
4.1 LCB-Guided Randomized Thresholding Algorithm
We present our main algorithm: LCB-Guided Randomized Thresholding (LGRT, Algorithm 1), that aims to learn the policy $\bm{\pi}^{b}$ that gives actions to agents whose incremental reward exceeds the threshold $\gamma$.
Algorithm 1 LCB-Guided Randomized Thresholding
1:  Input: $\bm{s},B,g(t),\phi^{m}(s,0),\gamma$
2:  Initialize: $\bm{s}_{0}=\bm{s}$, $\forall s\in\mathcal{S},\;m\in[M]:N_{s,m}^{(0)}=0$, $\hat{I}_{0}^{m}(s)=0$, $\hat{\phi}^{m}(s,1)=0$, $\mathrm{LCB}_{0}^{m}(s)=-\infty$
3:  for $t=0,1,\cdots,T$ do
4:     Retrieve $\bm{s}_{t}$, and obtain the set of agents in good states $\tilde{G}_{t}=\left\{m\colon\mathrm{LCB}_{t}^{m}({s}_{t}^{m})\geq\gamma\right\}$. Set $C=\tilde{G}_{t}$;
5:     Select $\max(0,B-|\tilde{G}_{t}|)$ agents uniformly from $[M]\setminus\tilde{G}_{t}$, and append them to $C$
6:     Uniformly select $B$ agents from $C$, and denote them by $\tilde{C}$. Set $a_{t+1}^{m}=1$ for $m\in\tilde{C}$, 00 otherwise
7:     Observe $(\bm{s}_{t+1},\bm{r}_{t+1})$
8:     Update cumulative incremental reward for each $m\in[M]$: $\hat{\phi}^{m}(s_{t}^{m},1)=\hat{\phi}^{m}(s_{t}^{m},1)+r^{m}_{t+1}\mathbbm{1}\left\{a_{t+1}^{m}=1\right\}$
9:     For each agent-state pair $(m,s)\in[M]\times\mathcal{S}$, update the lower confidence bound of the incremental rewards as follows:
10:         $N_{s,m}^{(t+1)}=N_{s,m}^{(t)}+\mathbbm{1}\left\{a_{t+1}^{m}=1\right\}\mathbbm{1}\left\{s=s_{t}^{m}\right\}$
11:         $\hat{I}_{t+1}^{m}(s)=\frac{\hat{\phi}^{m}(s,1)}{N_{s,m}^{(t+1)}}-\phi^{m}(s,0)$
12:         $\mathrm{LCB}_{t+1}^{m}(s)=\hat{I}_{t+1}^{m}(s)-\sqrt{\frac{g\left(N_{s,m}^{(t+1)}+2\right)}{N_{s,m}^{(t+1)}+2}}$
13:  end for
Algorithm 1 takes the initial state $\bm{s}$, budget $B$, threshold $\gamma$, and the state-action reward under no intervention as inputs. At each time step $t$, Algorithm 1 proceeds as follows: First, the algorithm obtains a set of agents, denoted by $\tilde{G}_{t}$, whose estimated lower confidence bounds of the incremental reward under the current state are above the threshold $\gamma$. If the cardinality of $\tilde{G}_{t}$ satisfies $|\tilde{G}_{t}|\geq B$, we uniformly sample $B$ agents from $\tilde{G}_{t}$ to give actions. Otherwise, we give all agents in $\tilde{G}_{t}$ actions and uniformly sample $B-|\tilde{G}_{t}|$ agents from the rest. After taking the action, we observe $\bm{s}_{t+1}$ and the corresponding reward $\bm{r}_{t+1}$. Next, we update the lower confidence bound $\mathrm{LCB}_{t+1}$. Let $\hat{\phi}^{m}(s,1)$ be the cumulative reward of each agent in each state, and $N_{s,m}^{(t+1)}$ be the number of times giving action 1 to agent $m$ when they are in state $s$. We update the estimated incremental reward $\hat{I}_{t+1}^{m}(s)$ for each agent and state using the empirical mean of reward under action 1. The $\mathrm{LCB}_{t+1}^{m}(s)$ is defined as $\hat{I}_{t+1}^{m}(s)-\sqrt{\frac{g\left(N_{s,m}^{(t+1)}+2\right)}{N_{s,m}^{(t+1)}+2}}$, where $g(t)=3\log(t)$.
Note that this lower confidence bound is only a function of $N_{s,m}^{(t)}$, the number of times action 1 is given to agent $m$ in state $s$. Compared to the typical anytime-valid LCB choice of $g(t)=\Theta(\log(t))$, our choice of $g$ is smaller, ensuring faster threshold crossing for agents with incremental reward above the threshold. We note that our algorithm shares similarities with Algorithm 1 in Michel et al. (2022), with a key distinction being the use of the LCB instead of the empirical mean when defining the candidate set $\tilde{G}_{t}$. This modification is crucial for achieving a superior regret bound compared to Michel et al. (2022). Intuitively, our choice of $g$ ensures that the LCB of an agent with an incremental reward smaller than the threshold will fall below the threshold more quickly. This design intuitively allows for faster elimination of “bad” agents, as demonstrated in the proof.
4.2 Theoretical Guarantee
Let $\Delta_{s_{\min}}^{m_{\min}}$ denote the smallest gap between the threshold and the incremental reward above the threshold (formally defined later). In this subsection, we establish that Algorithm 1 achieves a regret upper bound of $\mathcal{O}\left(\frac{1}{\Delta_{s_{\min}}^{m_{\min}}}\right)$ under Assumptions 4.1 and 4.2. This upper bound matches the lower bound $\Omega\left(\frac{1}{\Delta_{s_{\min}}^{m_{\min}}}\right)$up to some constants.
Assumption 4.1 ensures that all agents have strictly positive transition probabilities between states, even if these probabilities are arbitrarily small.
Assumption 4.1 (Positive Transition).
$P_{a}^{m}(s,s^{\prime})>0$ for all $s,s^{\prime}\in\mathcal{S}$, $a\in\mathcal{A}$, and $m\in[M]$.
The next assumption ensures the existence of a policy where there are always enough good arms at each time step.
Assumption 4.2 (Exists A Good Policy).
Assume that there exists a policy such that the cumulative cost is 0.
Assumption 4.2 can always be satisfied by appropriately choosing $\gamma$. For example, setting $\gamma=0$ guarantees the assumption holds. Before stating the main theorem (Theorem 4.4), we first introduce some notations. Let $\mathcal{G}$ be the set of good agent-state pairs, i.e., $\mathcal{G}:=\{(m,s):m\in[M],s\in\mathcal{S},I^{m}(s)\geq\gamma\}$. Let $\Delta_{s}^{m}:=\gamma-I^{m}(s)$. Let $\left(m_{\min},s_{\min}\right)$ be $\arg\min_{(m,s)\in\mathcal{G}}\left\{I^{m}(s)-\gamma\right\}$.
Next, Lemma 4.3 establishes that under Assumptions 4.1 and 4.2, two mild assumptions as discussed, we can ensure a strong guarantee that sufficient good agents exist for each possible state $\bm{s}\in\mathcal{S}^{M}$.
Lemma 4.3 (Sufficient Good Arms).
Under Assumptions 4.2 and 4.1, |$\left|\left\{(m,s^{m})\right\}\cap\mathcal{G}\right|\geq B$ for all $\bm{s}\in\mathcal{S}^{M}$.
Lemma 4.3 (proof in Appendix D.1) implies that although selecting a suboptimal agent for action may affect the future state distribution, this effect does not accumulate over time because sufficient good agents are always available. This ensures zero regret is achievable at each step. Next, we present our main theorem on the regret upper bound.
Theorem 4.4 (Regret Upperbound).
Assume $r_{t}^{m}$ is 1-subgaussian random variable for all $t\in[T]$. Let $g(t)=3\log(t)$. The regret, $\mathcal{R}(\pi)$ (Eq. (5)), satisfies:
$\mathcal{R}(\pi)=\min\left\{\mathcal{O}\left(\frac{|\mathcal{S}|M-|\mathcal{G}|}{\Delta_{s_{\min}}^{m_{\min}}}\right),\mathcal{O}\left(\sqrt{B|\mathcal{G}|(|\mathcal{S}|M-\mathcal{G})T}\right)\right\}.$ (8)
The specific constant will be formally stated in the proof (Appendix D). When the budget is one, Theorem 4.4 shows that our algorithm achieves a regret upper bound that is $\mathcal{O}\left(1/\Delta_{s_{\min}}^{m_{\min}}\right)$, which depends only on the gap between the incremental reward above the threshold and the threshold. This improves the regret bound of $\mathcal{O}\left(\max_{\Delta_{s}^{m}>0}1/\Delta_{s}^{m}\right)$ from Michel et al. (2022), which depends on the gap between incremental rewards below the threshold and the threshold—typically smaller than $\Delta_{s_{\min}}^{m_{\min}}$. Our result also aligns with the minimax lower bound established in Feng et al. (2024) (Theorem 3).
The optimality result in Theorem 4.4 relies on the presence of a sufficient number of good agents above the threshold. This contrasts with the findings of Feng et al. (2024), who demonstrate optimal regret regardless of the number of good arms above the threshold in the multi-armed bandit setting. Our problem, however, is fundamentally more challenging, making our analysis more difficult. Specifically, we address the contextual bandit setting where the context distribution evolves according to an MDP. In this setup, taking a suboptimal action has a long-term impact on future rewards by influencing the state distribution over subsequent periods. This complexity hinders a direct adaptation of the methods proposed by Feng et al. (2024) to our problem. While relaxing Assumption 4.2 may be possible, it falls outside the scope of this study and is left for future exploration.
Remark 4.5 (Sample Complexity).
While our algorithm shares a high-level idea with the online Whittle index algorithm (Wang et al., 2023b)—constructing confidence intervals for unknown statistics—the intervals in our model are naturally tighter. This is because, with more than two states, constructing confidence intervals for transition probabilities requires sufficient visits to state pairs $(s_{t},s_{t+1})$. In contrast, since our reward $\phi$ depends only on the current state $s_{t}$, our algorithm can gather information more efficiently.
Remark 4.6 (How to Select the Threshold $\gamma$?).
In our problem, there is an inherent trade-off between satisfying Assumption 4.2 and maximizing cumulative reward. While our reformulation focuses on the threshold identification problem, maximizing cumulative reward remains a desirable property in many real-world applications. Specifically, a smaller threshold $\gamma$ makes it easier to satisfy Assumption 4.2, as more state-agent pairs qualify as good. However, a low $\gamma$ also makes any arm above the threshold optimal, potentially leading to lower rewards. In the next section, we modify our algorithm, Algorithm 2, to greedily select actions based on the ordering of their confidence-bound estimators. This adjustment aims to enhance the cumulative reward achievable in practical applications.
5 Experiments
We numerically evaluate the performance of our algorithm, demonstrating that: 1) the regret is sublinear and converges to a constant under the budgeted thresholding contextual bandit reformulation , and 2) the achieved average cumulative reward, as defined in the finite-horizon restless bandit setting, outperforms that of existing online restless bandit algorithms. To comprehensively assess the performance of our algorithm for reward maximization across diverse scenarios, we conduct experiments under three settings:
  1. 1.
    Under Stochastic Dominance Assumption (Assumption 3.1): We test our modified algorithm, Algorithm 2, which is adapted from Algorithm 1 to better align with reward maximization objectives. This evaluation is conducted under the stochastic dominance assumption, a common framework in prior work (Wang et al., 2023b) and a key foundation for our optimality results.
  2. 2.
    Randomly Generated Instances: We evaluate our algorithm on randomly generated instances to demonstrate its robustness without relying on the stochastic dominance assumption. We modify both our algorithm (as reflected in Algorithm 3) and the benchmark algorithms to make them suitable for this setting and illustrate that our proposed algorithm outperforms existing modified benchmarks.
  3. 3.
    Adversarial Setting (No Action is Optimal): We test our algorithm in the setting where taking no action is optimal. Here, our algorithm performs comparably to existing benchmarks, further validating its adaptability. Results for this setting are deferred to Appendix E.2.
These results underscore the practicality of the proposed reformulation and algorithms in finite-horizon restless bandit problems.
Our modified algorithm, Algorithm 2, is designed to improve cumulative reward as well as the convergence rate under the budgeted thresholding contextual bandit reformulation. When sufficient good arms are identified, rather than randomly sampling from agents with LCB values above the threshold, we greedily select agents in $\tilde{G}_{t}$ in descending order of their LCB values. Further, when insufficient good arms are identified, we optimistically select agents based on the upper confidence bound (UCB) of the incremental reward rather than randomly sampling those with LCB below the threshold. The UCB at each time step is defined according to the anytime-valid confidence bound in Howard et al. (2021): $\mathrm{UCB}_{t+1}^{m}(s):=\hat{I}_{t+1}^{m}(s)+\sqrt{\frac{f(t,T)}{N_{s,m}^{(t+1)}+2}}$, where $f(t,T)=1.714^{2}\left(\log\log\left(N_{s,m}(t+1)+2\log\left(10T\right)\right)\right)$.
Algorithm 2 LCB-Guided Greedy Thresholding
1:  Input: $\bm{s},B,g(t),f(t),\phi^{m}(s,0),\gamma$
2:  Initialize: $\bm{s}_{0}=\bm{s}$, $\forall s\in\mathcal{S},\;m\in[M]:N_{s,m}^{(0)}=0$, $\hat{I}_{0}^{m}(s)=0$, $\hat{\phi}^{m}(s,1)=0$, $\mathrm{LCB}_{0}^{m}(s)=-\infty$, $\mathrm{UCB}_{0}^{m}(s)=\infty$
3:  for $t=0,1,\cdots,T$ do
4:     Retrieve $\bm{s}_{t}$, and obtain the set of agents in good states $\tilde{G}_{t}=\left\{m\colon\mathrm{LCB}_{t}^{m}({s}_{t}^{m})\geq\gamma\right\}$. Set $C=\tilde{G}_{t}$;
5:     Sort agents in $[M]\setminus\tilde{G}_{t}$ in descending order of $\text{UCB}_{t}^{m}(s_{t}^{m})$, select the first $\max(0,B-|\tilde{G}_{t}|)$ agents, and append them to $C$.
6:     Sort agents in $C$ in descending order of $\text{LCB}_{t}^{m}(s_{t}^{m})$, select the first $B$ agents, and denote them by $\tilde{C}$. Set $a_{t+1}^{m}=1$ for $m\in\tilde{C}$, 00 otherwise
7:     Observe $(\bm{s}_{t+1},\bm{r}_{t+1})$
8:     Update cumulative incremental reward for each $m\in[M]$: $\hat{\phi}^{m}(s_{t}^{m},1)=\hat{\phi}^{m}(s_{t}^{m},1)+r^{m}_{t+1}\mathbbm{1}\left\{a_{t+1}^{m}=1\right\}$
9:     For each agent-state pair $(m,s)\in[M]\times\mathcal{S}$, update the LCB and UCB of the incremental rewards as:
10:         $N_{s,m}^{(t+1)}=N_{s,m}^{(t)}+\mathbbm{1}\left\{a_{t+1}^{m}=1\right\}\mathbbm{1}\left\{s=s_{t}^{m}\right\}$
11:         $\hat{I}_{t+1}^{m}(s)=\frac{\hat{\phi}^{m}(s,1)}{N_{s,m}^{(t+1)}}-\phi^{m}(s,0)$
12:         $\mathrm{LCB}_{t+1}^{m}(s)=\hat{I}_{t+1}^{m}(s)-\sqrt{\frac{g\left(N_{s,m}^{(t+1)}+2\right)}{N_{s,m}^{(t+1)}+2}}$
13:        $\mathrm{UCB}_{t+1}^{m}(s)=\hat{I}_{t+1}^{m}(s)+\sqrt{\frac{f(t,T)}{N_{s,m}^{(t+1)}+2}}$
14:  end for
We will use this algorithm for Sections 5.1 and 5.2.
5.1 Regret Curve
Figure 1: Regret Curve as horizon $T$ grows.
To numerically verify Theorem 4.4, we set $M=2$, $B=1$, and $\mathcal{S}=\{0,1\}$. We set the expected reward of the state as $R^{m}(s)=s$. We randomly generated $10$ instances and chose three $\gamma$ values satisfying Assumption 4.2, labeling each $\gamma$ by the corresponding gap $\Delta_{s_{\min}}^{m_{\min}}$ between the threshold and the smallest incremental reward. The details of implementation are deferred to Appendix E.1.
For each $\gamma$, the regret was averaged over $100$ repetitions per instance (Figure 1). Figure 1 demonstrates sublinear regret for all choices of $\gamma$, with regret stabilizing at a constant as the horizon length increases, consistent with Theorem 4.4. Further, as $\Delta_{s_{\min}}^{m_{\min}}$ decreases, regret increases as it becomes harder to distinguish whether the incremental reward is above the threshold or not. This matches the results in Theorem 4.4.
5.2 Experiments With Stochastic Dominance
Next, we examine the average cumulative reward obtained by our algorithm in the online restless bandit setting (Section 2.2) under stochastic dominance assumption, and compare it with the following benchmarks: 1) WIQL (Biswas et al., 2021), 2) variants of UCWhittle (UCWhittle-UCB-Privileged, UCWhittlePv-Privileged) (Wang et al., 2023b), 3) Oracle Whittle (Whittle, 1988), 4) Oracle Greedy and 5) Random Policy, which aligns with the experimental setting of Wang et al. (2023b). We emphasize that our experiments involve heterogeneous agents, a setting in which no algorithm has been established as optimal. To evaluate the relative quality of the policies to which the online algorithms converge, we incorporate two benchmark algorithms with full MDP knowledge: Oracle Whittle and Oracle Greedy. We note that in this setting, neither algorithm is guaranteed to be optimal.
  • Oracle Whittle: A variant of Whittle’s index policy with access to the ground truth transitions and rewards, optimized for maximizing discounted reward (as described in Appendix A).
  • Oracle Greedy: The greedy policy we defined (Algorithm 2) with knowledge of the ground truth incremental reward.
  • WIQL (Biswas et al., 2021): A Q-learning based approach to learn the value function of each arm at each state by interacting with the RMAB instance. We use the same learning rate ($\alpha$) as in Biswas et al. (2021).
  • UCWhittle (Wang et al., 2023b): It constructs the confidence interval of the unknown transition probability and then chooses one to calculate Whittle’s index policy. We compare with two different choices.
    • UCWhittle-UCB (Wang et al., 2023b): It uses the upper confidence bounds of transition probabilities.
    • UCWhittle-penalty (Wang et al., 2023b): It constructs a confidence interval and chooses the transition probability that will maximize $V_{\lambda}^{m}(s)$ (Eq. (A.13)) for some adaptively updated $\lambda$.
In all of the settings, the discount factor $\beta$ was set to $0.9$. We note that though $\beta=1$ was chosen in Biswas et al. (2021), $\beta=0.9$ has better empirical performance, which aligns with the choice in Wang et al. (2023b). [^3] Code is available at github. Our results differ from those of Wang et al. (2023b) due to detected bugs in their code; a detailed discussion is provided in Appendix E.2.
Since our algorithm has access to $\phi^{m}(s,0)$, for fairness, we assume that UCWhittle algorithms have full knowledge of the transition probability under no intervention. This is a stronger assumption, as accurately estimating transition probabilities would require significantly more offline data than estimating $\phi^{m}(s,0)$. Since it is challenging to determine how Q-learning algorithms could effectively use additional information from the no-intervention setting, we did not provide WIQL with any extra information.
We set $M=50$, $\mathcal{S}=\{0,1\}$, and compare the results under three different budget settings: $B=5,10,20$. The expected reward is set to be $R^{m}(s)=s,\forall m\in[M],s\in\mathcal{S}$ and randomly generate $50$ instances with different transition matrices.
Figure 2: Average cumulative reward under the stochastic dominance assumption, evaluated across three budget settings: $B=5$ (left column), $10$ (middle column), $20$ (right column). Results are averaged over $50$ instance satisfying Assumption 3.1, each with $13$ repetitions. Top row: noiseless reward. Bottom row: noisy reward.
To enforce the stochastic dominance assumption, we modify randomly generated instances by checking if there exist $s,m$ such that $P_{1}^{m}(s,1)<P_{0}^{m}(s,1)$. If so, we swap these transitions to satisfy Assumption 3.1. In Figure 2, each line is averaged over $50$ instances, each with $13$ repetitions. We tested our algorithm, LGGT, with two threshold values $0.1$ and $0.4$. We consider two reward settings: 1) a noiseless case where $r_{t}\in\{0,1\}$, and 2) a noisy setting where the rewards at state $s\in\mathcal{S}$ follow a normal distribution with mean $R^{m}(s)$ and variance 1.
Under the noiseless reward setting (Figure 2 top row), in all three budget settings, LGGT with $\gamma=0.1$ consistently outperforms online restless bandit algorithms within a horizon of $1000$. Additionally, as the budget increases, the performance gap between LGGT ($\gamma=0.1$) and WIQL widens. At first glance, this may seem counterintuitive—since lower budgets reduce action frequency, one might expect learning the Q-function in WIQL to be harder. Indeed, the smallest performance gap occurs at the lowest budget, highlighting the difficulty of learning in low-budget settings. In resource-limited settings, while our algorithm is more efficient in terms of learning, its potentially larger suboptimality gap compared to benchmark algorithms may negatively impact its overall finite-sample performance. As shown in Figure 2 (top left), LGGT quickly converges to a stable yet suboptimal policy, whereas WIQL converges more slowly. However, given a sufficiently long horizon, WIQL will eventually converge to the Oracle Whittle policy. Nevertheless, LGGT ($\gamma=0.1$) outperforms WIQL when the horizon is short even when the budget is low, highlighting the practicality of our algorithm in the online restless bandit settings. Similarly, we observe that UCWhittleUCBPriv converges slowly, and it is worse than LGGT ($\gamma=0.1$) under all three budget settings. Further, UCWhittlePvPriv performs the worst when the number of agents is large.
In the noisy reward setting (Figure 2, bottom row), we excluded UCWhittle policies since constructing Whittle’s index requires precise reward knowledge. Handling noise would require carefully calibrated confidence intervals for both rewards and transitions, which is beyond the scope of this work. Additionally, since UCWhittleUCBPriv underperformed WIQL in the noiseless setting, we did not include the UCWhittle family in the noisy setting. In all three budget settings, LGGT with $\gamma=0.1$ consistently outperforms WIQL. Moreover, the gap between LGGT and WIQL widens when compared to noiseless setting. This indicates that LGGT adapts better to the noise in rewards, demonstrating the robustness of our algorithm.
Lastly, we compare the performance of Oracle Whittle and Oracle Greedy. With the stochastic dominance assumption, Oracle Whittle consistently outperforms Oracle Greedy across all budget settings, with the performance gap narrowing as the budget increases. Despite this gap, LGGT outperforms all online restless bandit benchmarks, demonstrating the benefit of rapid convergence in finite-horizon scenarios.
5.3 Experiments Without Stochastic Dominance
Figure 3: Average cumulative reward without the stochastic dominance assumption, evaluated across three budget settings: $B=5$ (left column), $10$ (middle column), $20$ (right column). Results are averaged over $50$ instances, each with $13$ repetitions. Top row: noiseless reward. Bottom row: noisy reward.
In this subsection, we evaluate the performance of all algorithms in the absence of the stochastic dominance assumption. As discussed in Remark 3.5, when stochastic dominance does not hold, we might favor not using the entire budget at each time step. Therefore, we adapt all algorithms to accommodate scenarios where the stochastic dominance assumption may not hold.
To adapt LGGT, we restrict interventions to agents whose incremental reward is estimated to be positive, as outlined in Algorithm 3. In particular, when the incremental reward of an agent-state pair $(m,s^{m})$ is negative, no action will be taken by agent $m$ when they are in state $s^{m}$. We will denote this algorithm as LCB-Guided Greedy Thresholding With UCB Filtering (LGGTWUF). Similarly, for policies based on Whittle’s index, we restrict interventions to agents with positive index values, i.e., when Whittle’s index of an agent-state pair $(m,s^{m})$ is negative, no action will be taken by agent $m$ when they are in state $s^{m}$. We note that the optimality of these adaptations in the absence of stochastic dominance has not been established.
Algorithm 3 LCB-Guided Greedy Thresholding With UCB Filtering
1:  Input: $\bm{s},B,g(t),f(t),\phi^{m}(s,0),\gamma$
2:  Initialize: $\bm{s}_{0}=\bm{s}$, $\forall s\in\mathcal{S},\;m\in[M]:N_{s,m}^{(0)}=0$, $\hat{I}_{0}^{m}(s)=0$, $\hat{\phi}^{m}(s,1)=0$, $\mathrm{LCB}_{0}^{m}(s)=-\infty$, $\mathrm{UCB}_{0}^{m}(s)=\infty$, $\mathrm{TEST}_{0}^{m}(s)=\infty$
3:  for $t=0,1,\cdots,T$ do
4:     Retrieve $\bm{s}_{t}$, and obtain the set of agents in good states $\tilde{G}_{t}=\left\{m\colon\mathrm{LCB}_{t}^{m}({s}_{t}^{m})\geq\gamma\right\}$. Set $C=\tilde{G}_{t}$;
5:     Sort agents in $[M]\setminus\tilde{G}_{t}$ in descending order of $\text{UCB}_{t}^{m}(s_{t}^{m})$, select the first $\max(0,B-|\tilde{G}_{t}|)$ agents, and append them to $C$.
6:     Sort agents in $C$ in descending order of $\text{LCB}_{t}^{m}(s_{t}^{m})$, select the first $B$ agents, and denote them by $\tilde{C}$.
7:     Obtain the set of agents whose incremental reward is positive $\tilde{L}_{t}:=\{m:\mathrm{TEST}_{t}^{m}(s_{t}^{m})\geq 0\}$. Set $a_{t+1}^{m}=1$ for $m\in\tilde{C}\cap\tilde{L}_{t}$, 00 otherwise
8:     Observe $(\bm{s}_{t+1},\bm{r}_{t+1})$
9:     Update cumulative incremental reward for each $m\in[M]$: $\hat{\phi}^{m}(s_{t}^{m},1)=\hat{\phi}^{m}(s_{t}^{m},1)+r^{m}_{t+1}\mathbbm{1}\left\{a_{t+1}^{m}=1\right\}$
10:     For each agent-state pair $(m,s)\in[M]\times\mathcal{S}$, update the LCB and UCB of the incremental rewards as:
11:         $N_{s,m}^{(t+1)}=N_{s,m}^{(t)}+\mathbbm{1}\left\{a_{t+1}^{m}=1\right\}\mathbbm{1}\left\{s=s_{t}^{m}\right\}$
12:         $\hat{I}_{t+1}^{m}(s)=\frac{\hat{\phi}^{m}(s,1)}{N_{s,m}^{(t+1)}}-\phi^{m}(s,0)$
13:         $\mathrm{LCB}_{t+1}^{m}(s)=\hat{I}_{t+1}^{m}(s)-\sqrt{\frac{g\left(N_{s,m}^{(t+1)}+2\right)}{N_{s,m}^{(t+1)}+2}}$
14:         $\mathrm{TEST}_{t+1}^{m}(s)=\hat{I}_{t+1}^{m}(s)+\sqrt{\frac{g\left(N_{s,m}^{(t+1)}+2\right)}{N_{s,m}^{(t+1)}+2}}$
15:        $\mathrm{UCB}_{t+1}^{m}(s)=\hat{I}_{t+1}^{m}(s)+\sqrt{\frac{f(t,T)}{N_{s,m}^{(t+1)}+2}}$
16:  end for
Similar to the case where stochastic dominance holds, we tested our algorithm, LGGTWUF, with two threshold values $\gamma=-0.9$ and 00. Note that when stochastic dominance does not hold, the incremental reward may become negative. Consequently, the cases with $\gamma=-0.9$ and 00 exhibit similar behavior to those with $\gamma=0.1$ and $0.4$ under stochastic dominance. We consider both noiseless and noisy settings and include Oracle Whittle and Oracle Greedy as benchmarks.
In the noiseless reward setting (Figure 3 top row), LGGT with $\gamma=-0.9$ consistently outperforms online restless bandit algorithms across three budget settings within a horizon of $1000$. This phenomenon is consistent with the results under the stochastic dominance assumption, demonstrating that the strong performance of our proposed algorithm extends beyond scenarios where stochastic dominance holds. Furthermore, UCWhittlePvPriv performs close to RandomPolicy when the stochastic dominance assumption does not hold and when the number of agents is large.
In the noisy reward setting (Figure 3 bottom row), LGGT with $\gamma=-0.9$ still consistently outperforms online restless bandit algorithms across all budget settings within a horizon of $1000$. However, the performance gap between LGGT and WIQL shrinks. We attribute this to the use of a naive conservative test (upper confidence bound) in LGGT to determine whether the incremental reward is positive, which limits its performance in the noisy setting. We hypothesize that replacing this test with more sophisticated procedures could potentially improve the convergence rate.
Finally, when stochastic dominance does not hold, the performance gap between the Oracle Whittle and Oracle Greedy is smaller than in Section 5.2. This highlights the practical utility of our algorithm in general online restless bandit settings.
6 Conclusion
To the best of our knowledge, this is the first paper to establish the optimality of a greedy policy in finite-horizon restless bandits. Based on this, we reformulate the online restless bandit as a budgeted thresholding contextual bandit and provide an algorithm with a minimax constant regret bound. Our algorithm builds on the intuition that a policy converging faster to a suboptimal solution can outperform those converging more slowly to a more optimal one in finite-sample settings. We further demonstrate the superior performance of our algorithm through numerical experiments.
Appendix A Whittle’s Index
Whittle’s index is computed for each agent. At each time $t$, given $\bm{s}_{t}$, the top $B$ agents with the highest indices are selected. Whittle’s index policy has two variants, each aimed at maximizing a different objective in an infinite horizon setting: one for maximizing average reward and the other for maximizing discounted reward.
Average Reward: In this case, Whittle’s index policy is designed to approximately find the policy $\bm{\pi}_{d}$ that maximizes the average reward,
$\max_{\bm{\pi}_{d}}~{}$ $\liminf_{T\to\infty}\frac{1}{TM}\mathbb{E}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]$
s.t. $\sum_{m=1}^{M}a_{t}^{m}\leq B,\forall t\in[T].$ (A.9)
We will denote $J^{*}$ to be the optimal value of Problem A.9. Instead of directly solving Problem A.9, Whittle deals with the Lagrangian relaxation of Problem A.9, which is stated in Problem A.10
$\max_{\bm{\pi}_{d}}~{}\liminf_{T\to\infty}\frac{1}{TM}\mathbb{E}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)+\lambda\left(1-a_{t+1}^{m}\right)\mid\bm{s}_{0}\right].$ (A.10)
The key observation to derive Whittle’s index policy is that Problem A.10 can be solved by addressing the following dynamic programming separately for each agent
$V^{m}_{\lambda}(s)$ $=\max_{a\in\{0,1\}}Q^{m}(s,a)$
$Q^{m}_{\lambda}(s,a)+\beta^{*}_{m}$ $=R^{m}(s)-\lambda\mathbbm{1}\{a=1\}$
$+\sum_{s^{\prime}\in\mathcal{S}}P^{m}_{a}(s,s^{\prime})V^{m}(s^{\prime}),$
where $\beta^{*}_{m}$ is the maximized average reward for arm $m$ without any constraint, i.e.
$\beta^{*}_{m}:=\max~{}\liminf_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid s_{0}^{m}\right].$
The Whittle’s index $\lambda^{m}_{s}$ is defined as the multiplier $\lambda_{s}^{m}$ when it is equally favorable between choosing action $1$ and action 00, i.e. $Q^{m}_{\lambda^{m}_{s}}(s,1)=Q^{m}_{\lambda^{m}_{s}}(s,0)$. The Whittle’s index policy would then give action to $B$ agents with the highest Whittle’s index at each time step $t$. Denote this as $\bm{\pi}^{\text{Whittle}}$ and the average reward of Whittle’s index policy as $J^{\text{Whittle}}$. The performance of Whittle’s index policy is characterized in Theorem A.1.
Theorem A.1 (Theorem 2, Weber & Weiss 1990).
Suppose all arms are homogeneous, i.e., $P_{a}^{m}=P_{a}^{m^{\prime}},R^{m}=R^{m^{\prime}},\forall m\neq m^{\prime}$, then the Whittle’s index policy is asymptotically optimal,
$\lim_{M\to\infty}J^{\text{Whittle}}=\lim_{M\to\infty}J^{*}.$
Discounted Reward: In this case, Whittle’s index policy is designed to approximately find the policy $\bm{\pi}_{d}$ that maximizes the discounted reward for a fixed discount factor $\beta$,
$\max_{\bm{\pi}_{d}}~{}$ $\sum_{t=0}^{\infty}\frac{1}{M}\mathbb{E}\left[\sum_{m=1}^{M}\beta^{t}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]$
s.t. $\sum_{m=1}^{M}a_{t}^{m}\leq B,\forall t\in[T].$ (A.11)
Similarly, instead of directly solving Problem A.11, Whittle deals with the Lagrangian relaxation of Problem A.11, which is stated in Problem A.12
$\max_{\bm{\pi}_{d}}~{}\sum_{t=0}^{\infty}\frac{1}{M}\mathbb{E}\left[\sum_{m=1}^{M}\beta^{t}\left(R^{m}\left(s_{t}^{m}\right)+\lambda\left(1-a_{t+1}^{m}\right)\right)\mid\bm{s}_{0}\right].$ (A.12)
Again, the same key observation is that Problem A.12 can be solved by addressing the following dynamic programming separately for each agent
$V^{m}_{\lambda}(s)$ $=\max_{a\in\{0,1\}}Q^{m}(s,a)$
$Q^{m}_{\lambda}(s,a)$ $=R^{m}(s)-\lambda\mathbbm{1}\{a=1\}$
$+\beta\sum_{s^{\prime}\in\mathcal{S}}P^{m}_{a}(s,s^{\prime})V^{m}(s^{\prime}).$ (A.13)
The definition of Whittle’s index will be the same in the average reward case.
A.1 Generalized Regret in Online Restless Bandits
Given $K$ episodes each with a horizon length of $T$ and the set of online Whittle’s index policy at each episode $k\in[K]$, $\bm{\omega}:=\{\bm{\omega}_{k}\}_{k\in[K]}$, the regret $\mathcal{R}_{f}(\bm{\omega})$:
$\mathcal{R}_{f}(\bm{\omega}):$ $=\frac{1}{K}\sum_{k=1}^{K}\Bigg{(}\mathbb{E}_{\bm{\pi}^{\text{Whittle}}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]-\mathbb{E}_{\bm{\omega}_{k}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]\Bigg{)}.$ (A.14)
We note that our problem setting corresponds to the scenario where $K=1$.
Existing regret bounds on online Whittle’s index policy are only meaningful when the number of episodes is sufficiently large. To illustrate this, Wang et al. (2023b) shows that their algorithm has a regret bound of the order $\mathcal{O}\left(T\sqrt{TK}\right)$. If we only have $1$ episode, this will be equivalent to saying that the regret bound is $\mathcal{O}(T^{3/2})$. The hardness lies in estimating the MDP: Azar et al. 2017 show that when $B=M=1$, the lower bound of the regret is $\Omega(T\sqrt{K})$. This lower bound indicates that when directly dealing with an MDP, it is impossible to get a regret bound that is sublinear in $T$.
Appendix B Counter Example for Heterogeneous Agents
In this section, we will construct an instance involving heterogeneous agents where the greedy policy is not optimal. Say that each agent is associated with two state: State 0 and State 1, and the agent can take either Action $1$ (intervention) or Action 00 (no intervention). Assume for now that there are 4 agents with different transition probabilities. The first two agents fall under Type 1: their transition probability under Action 1 is shown in Figure B.4 and their transition probability under Action 0 is shown in Figure B.5. The other two agents fall under Type 2: their transition probability under Action 1 is shown in Figure B.6 and their transition probability under Action 0 is shown in Figure B.7.
[…]
Figure B.4: Dynamic of Agent Type 1 with Intervention
[…]
Figure B.5: Dynamic of Agent Type 1 without Intervention
[…]
Figure B.6: Dynamic of Agent Type 2 with Intervention
[…]
Figure B.7: Dynamic of Agent Type 2 without Intervention
Suppose the budget equals one, meaning that we can only give one agent intervention at each time. We observe that under our problem setup Type 2 agents always have a higher incremental reward $\phi^{2}(1,1):=\frac{3}{4}\cdot 1,\phi^{2}(0,1):=\frac{3}{4}\cdot 1,\phi^{2}(s,0)=0,I^{2}(1)=I^{2}(0)=\frac{3}{4}$, opposed to $\phi^{1}(0,1)=\frac{2}{3}\cdot 1,\phi^{1}(1,1)=\frac{1}{2}\cdot 1,\phi^{1}(0,0)=0,\phi^{1}(1,0)=\frac{1}{3},I(1)=\frac{2}{3},I(0)=\frac{1}{6}$. However, this does not imply that the optimal policy is to always take Action 1 for Type 2 agents. If all resources are allocated to Type 2 agents, the expected reward is $3/4$ per time step summed across all agents. Alternatively, if resources are devoted to Type 1 agents, alternating actions between the two Type 1 agents, the expected reward is $2/3+2/3\times 1/3=8/9$ per Type 1 agent over two time periods (where an action is taken in one period but not in the other). To illustrate this point, we ran the following experiment: we set the horizon to be $20$ and repeated the simulation $300$ times. We compare the performance of the following two policies in terms of cumulative rewards: Policy 1) choosing one of the Type 2 agents for intervention at each time, and Policy 2) alternating intervention between the two Type 1 agents. The observed cumulative for Policies 1 and 2 are 15.02 and 16.78, respectively. This shows that the greedy policy is not optimal under all heterogeneous agents cases even under Assumption 3.1.
Appendix C Technical Details In Section 3
First, recall that at each time $t$ of an offline restless bandit problem, the learner observes the set of states of $M$ agents, denoted by $\bm{s}_{t}:=\left(s_{t}^{1},s_{t}^{2},\cdots,s_{t}^{M}\right)$, and receives the corresponding reward $\bm{r}_{t}:=(r_{t}^{1},\cdots,r_{t}^{M})$. Based on the observed states, the learner determines the action $\bm{a}_{t+1}:=\left(a_{t+1}^{1},a_{t+1}^{2},\cdots,a_{t+1}^{M}\right)$, according to a deterministic policy $\pi^{d}_{t+1}(\bm{s}):\mathcal{S}^{M}\rightarrow\{0,1\}^{M}$. The learner is provided with a budget constraint, allowing at most $B$ arms to be pulled at each time step, i.e., $\sum_{m\in[M]}a_{t+1}^{m}\leq B$. After pulling up to $B$ arms, the learner observes the next state $\bm{s}_{t+1}$, generated according to $s_{t+1}^{m}\sim P_{a_{t+1}^{m}}^{m}\left(s_{t}^{m},s_{t+1}^{m}\right)$, and receives a feedback $\bm{r}_{t+1}$.
To ease notation, in Section C we will use $q^{a}_{ss^{\prime}}(m)$ to denote $P_{a}^{m}(s,s^{\prime})$. Before presenting the proofs, we reformulate the online restless bandit problem as a single large MDP, where the state and action are defined as tuples comprising the composite states and actions of all agents. This reformulation is valid because the agents/arms in the restless bandit problem are independent. This MDP can be described as $\left(\mathcal{S}^{M},\mathcal{F},q,\bm{R}\right)$, where $\mathcal{F}$ is defined as
[…]
Let $q_{\bm{s}\bm{s}^{\prime}}^{\bm{a}}:=\mathbb{P}(\bm{s}^{\prime}|\bm{s},\bm{a})=P_{\bm{a}}(\bm{s},\bm{s}^{\prime})=\prod_{m=1}^{M}q^{a}_{ss^{\prime}}(m)$, and let $\bm{R}(\bm{s})=\sum_{m=1}^{M}r^{m}\left(s^{m}\right)$. Then for each policy $\bm{\pi}$ and each time-step $h$, we can define the value function $V_{h}^{\bm{\pi}}(\bm{s})$ as
$\mathbb{E}_{\bm{\pi}}\left[\sum_{t=h}^{T}\bm{R}(\bm{s}_{t})\mid\bm{s}_{h}=\bm{s}\right].$ (C.15)
With this definition, Problem (1) is equivalent to
$\max_{\bm{\pi}}V_{0}^{\bm{\pi}}\left(\bm{s}_{0}\right).$
Again, we use $\bm{\pi}^{*}:=\{\pi_{t}^{*}\}_{t\in[T]}$ to denote the maximizer, and let $V_{h}^{*}$ be the value function under policy $\bm{\pi}^{*}$. We will use this notation in Appendix C.2, C.3, C.4, and D.
Next, we state the proofs.
C.1 Proof of Lemma 3.2
See 3.2
Proof.
$\Rightarrow$: When Assumption 3.1 holds, suppose there exists $s\in\mathcal{S}$ such that $q_{s1}^{1}<q_{s1}^{0}$. Without loss of generality (WLOG), assume $s=1$. We have for any increasing function $f$:
$\mathbb{E}_{s^{\prime}\sim q_{1s^{\prime}}^{1}}\left[f(s^{\prime})\right]-\mathbb{E}_{s^{\prime}\sim q_{1s^{\prime}}^{0}}\left[f(s^{\prime})\right]$ $=f(1)q^{1}_{11}+f(0)q^{1}_{10}-f(1)q^{0}_{11}-f(0)q^{0}_{10}$
$=f(1)\left(q^{1}_{11}-q^{0}_{11}\right)+f(0)\left(q^{1}_{10}-q^{0}_{10}\right)$
$=\left(f(1)-f(0)\right)\left(q^{1}_{11}-q^{0}_{11}\right)$
$<0,$
where the last inequality holds because $f$ is increasing and the assumption that $q^{1}_{11}<q^{0}_{11}$. This will contradict Assumption 3.1. Therefore, we must have for all $s\in\mathcal{S}$, $q_{s1}^{1}\geq q_{s1}^{0}$.
:absent:$\Leftarrow$:⇐ : When for all $s\in\mathcal{S}$, $q_{s1}^{1}\geq q_{s1}^{0}$, we will prove Eq. (7) separately for $s=0$ and $s=1$. When $s=1$, we have
$\mathbb{E}_{s^{\prime}\sim q_{1s^{\prime}}^{1}}\left[f(s^{\prime})\right]-\mathbb{E}_{s^{\prime}\sim q_{1s^{\prime}}^{0}}\left[f(s^{\prime})\right]$ $=f(1)q^{1}_{11}+f(0)q^{1}_{10}-f(1)q^{0}_{11}-f(0)q^{0}_{10}$
$=f(1)\left(q^{1}_{11}-q^{0}_{11}\right)+f(0)\left(q^{1}_{10}-q^{0}_{10}\right)$
$=\left(f(1)-f(0)\right)\left(q^{1}_{11}-q^{0}_{11}\right)$
$\geq 0,$
where the last inequality holds because $f$ is increasing and the assumption that $q^{1}_{11}\geq q^{0}_{11}$. Similarly, we can show that $\mathbb{E}_{s^{\prime}\sim q_{0s^{\prime}}^{1}}\left[f(s^{\prime})\right]\geq\mathbb{E}_{s^{\prime}\sim q_{0s^{\prime}}^{0}}\left[f(s^{\prime})\right]$. This completes the proof. ∎
C.2 Proof of Lemma 3.3
See 3.3
Proof.
We will use Bellman optimality equation to prove the theorem. Specifically, we only need to prove that for any $h$, the maximizer $a$ of the following problem,
$V_{h}^{*}(s)=R(s)+\max_{a}\sum_{s^{\prime}}q_{ss^{\prime}}^{a}V_{h+1}^{*}(s),$
aligns with $\pi_{h+1}^{\text{greedy}}(s)$. With Assumption 3.1, this is equivalent to showing $V_{h+1}^{*}(s)$ is an increasing function in $s$. To do this, we will discuss two cases: 1) $q_{11}^{1}\geq q_{01}^{1}$ and 2) $q_{11}^{1}\leq q_{01}^{1}$.
Case 1:
$q^{1}_{11}\geq q^{1}_{01}$. We will use backward induction to prove the following statements:
$\forall t\leq T,~{}$ $V^{*}_{t}(1)\geq V_{t}^{*}(0);$ (C.16)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.17)
When $t=T$, the statements hold trivially. Suppose the statements holds for all $t\geq h+1$, when $t=h$, we will first prove that the greedy policy is optimal assuming $V^{*}_{h+1}(1)\geq V_{h+1}^{*}(0)$. By Bellman optimality equation, we have
$V_{h}^{*}(s)=R(s)+\max_{a}\sum_{s^{\prime}}q^{a}_{ss^{\prime}}V_{h+1}^{*}\left(s^{\prime}\right).$
Since $V_{h+1}^{*}$ is increasing function, by Assumption 3.1, we have
$\sum_{s^{\prime}}q^{1}_{ss^{\prime}}V_{h+1}^{*}\left(s^{\prime}\right)\geq\sum_{s^{\prime}}q^{0}_{ss^{\prime}}V_{h+1}^{*}\left(s^{\prime}\right).$
This implies that action 1 is the maximizer, and the optimal policy will choose to give intervention at $t$, indicating Eq. (C.17) holds. Next, assuming that $\pi_{h+1}^{\text{greedy}}=\pi_{h+1}^{*}$, we prove for $V^{*}_{h}(1)\geq V_{h}^{*}(0)$:
$V_{h}^{*}(1)-V_{h}^{*}(0)$ $=R(1)-R(0)+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1,s^{\prime}}V_{h+1}^{*}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0,s^{\prime}}V_{h+1}^{*}(s^{\prime})$
$\geq\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}(1)-V_{h+1}^{*}(0)\right)$
$\geq 0.$
Case 2:
$q^{1}_{11}\leq q^{1}_{01}$. Similarly, we use backward induction to prove the following statements:
$\forall t\leq T,~{}$ $0\leq V^{*}_{t}(1)-V_{t}^{*}(0)\leq R(1)-R(0);$ (C.18)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.19)
This trivially holds for $t=T$. Suppose this holds for $t\geq h+1$, when $t=h$, we can prove Eq. (C.19) for $t=h$ using exactly the same logic. Assuming that $\pi_{h+1}^{\text{greedy}}=\pi_{h+1}^{*}$ and that Eq. (C.18) hold for $t=h+1$, we have
$V_{h}^{*}(1)-V_{h}^{*}(0)$ $=R(1)-R(0)+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1,s^{\prime}}V_{h+1}^{*}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0,s^{\prime}}V_{h+1}^{*}(s^{\prime})$
$=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}(1)-V_{h+1}^{*}(0)\right)$
$\leq R(1)-R(0).$
This will give us the RHS of Eq. (C.18) for $t=h$. Next, we will show the LHS of Eq. (C.18) for $t=h$ under the same assumption that $\pi_{h+1}^{\text{greedy}}=\pi_{h+1}^{*}$ and that Eq. (C.18) hold for $t=h+1$.
$V_{h}^{*}(1)-V_{h}^{*}(0)$ $=R(1)-R(0)+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1,s^{\prime}}V_{h+1}^{*}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0,s^{\prime}}V_{h+1}^{*}(s^{\prime})$
$=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}(1)-V_{h+1}^{*}(0)\right)$
$\geq R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(R(1)-R(0)\right)$
$=\left(q^{1}_{11}+q_{00}^{1}\right)\left(R(1)-R(0)\right)$
$\geq 0.$
Together, they will give us Eq. (C.18) for $t=h$. This completes the proof. ∎
C.3 Proof of Lemma 3.4
See 3.4
Proof.
WLOG, assume $M=2,B=1$. Again, using Bellman optimal equation, we only need to prove that for any $t\leq T$, the following two equations hold for any $\left(s_{t}^{1},s_{t}^{2}\right)\in\mathcal{S}^{2}$:
$\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{1}_{s_{t}^{1},s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right)$
$\geq\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{t}^{1}s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right)$ (C.20)
$\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{t}^{1}s_{t+1}^{1}}(1)q^{1}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right)$
$\geq\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{t}^{1}s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right).$ (C.21)
It is easy to see that these two equations are symmetric, thus, we will only provide the proof for the first inequality. Same as the proof of Lemma 3.3, we will prove the result under two cases: 1) $q^{1}_{11}(1)\geq q^{1}_{01}(1)$ and 2) $q^{1}_{11}(1)\leq q^{1}_{01}(1)$.
Case 1:
$q^{1}_{11}(1)\geq q^{1}_{01}(1)$. We will use backward induction to prove the following statements:
$\forall t\leq T,~{}$ $f_{s,t}(1)\geq f_{s,t}(0),\forall s\in\mathcal{S};$ (C.22)
$\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{1}_{s_{t}^{1},s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right)$
$\geq\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{t}^{1}s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right),\forall s_{t}^{i}\in\mathcal{S},$ (C.23)
where $f_{s,t}\left(s^{\prime}\right):=\sum_{s_{t+1}^{2}\in\mathcal{S}}q^{0}_{ss_{t+1}^{2}}\left(2\right)V_{t+1}^{*}\left(\left(s^{\prime},s_{t+1}^{2}\right)\right)$. They hold trivially when $t=T$. Suppose the statements hold for $t\geq h+1$, when $t=h$, we will first prove Eq. (C.22) for $t=h$ assuming that Eq.(C.22) holds for $t=h+1$.
$f_{s,h}(1)-f_{s,h}(0)$ $=\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{t+1}^{2}}\left(2\right)V_{h+1}^{*}\left(\left(1,s_{h+1}^{2}\right)\right)-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{t+1}^{2}}\left(2\right)V_{h+1}^{*}\left(\left(0,s_{h+1}^{2}\right)\right)$
$=\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(1)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(0)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$=R^{1}(1)-R^{1}(0)+\left(q^{1}_{11}(1)-q^{1}_{01}(1)\right)\left(\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(f_{s_{h+1}^{2},h+1}(1)-f_{s_{h+1}^{2},h+1}(0)\right)\right)$
$\geq 0.$
Next, we want to show Eq. (C.23) for $t=h$ assuming that Eq. (C.22) holds for $t=h$.
$\sum_{\left(s_{h+1}^{1},s_{h+1}^{2}\right)\in\mathcal{S}^{2}}q^{1}_{s_{h}^{1}s_{h+1}^{1}}(1)q^{0}_{s_{h}^{2}s_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(s_{h+1}^{1},s_{h+1}^{2}\right)\right)$
$-\sum_{\left(s_{h+1}^{1},s_{h+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{h}^{1}s_{h+1}^{1}}(1)q^{0}_{s_{h}^{2}s_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(s_{h+1}^{1},s_{h+1}^{2}\right)\right)$
$=\sum_{s_{h+1}^{1}}q^{1}_{s_{h}^{1}s_{h+1}^{1}}(1)f_{s_{h}^{2},h}\left(s_{h+1}^{1}\right)-\sum_{s_{h+1}^{1}}q^{0}_{s_{h}^{1}s_{h+1}^{1}}(1)f_{s_{h}^{2},h}\left(s_{h+1}^{1}\right)$
$\geq 0,$
where the last inequality holds because of Eq. (C.22) and Assumption 3.1.
Case 2:
$q^{1}_{11}(1)\leq q^{1}_{01}(1)$. We will use the backward induction to prove the following statements:
$\forall t\leq T,~{}$ $0\leq f_{s,t}(1)-f_{s,t}(0)\leq R^{1}(1)-R^{1}(0),\forall s\in\mathcal{S};$ (C.24)
$\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{1}_{s_{t}^{1},s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right)$
$\geq\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{t}^{1}s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right),\forall s_{t}^{i}\in\mathcal{S}.$ (C.25)
They trivially hold for $t=T$. Suppose the statements hold for $t\geq h+1$, when $t=h$, we will first prove the RHS of Eq. (C.24) for $t=h$ assuming that Eq. (C.24) holds for $t=h+1$.
$\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(1,s_{h+1}^{2}\right)\right)-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(0,s_{h+1}^{2}\right)\right)$
$=\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(1)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(0)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$=R^{1}(1)-R^{1}(0)+\left(q^{1}_{11}(1)-q^{1}_{01}(1)\right)\left(\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(f_{s_{h+1}^{2},h+1}(1)-f_{s_{h+1}^{2},h+1}(0)\right)\right)$
$\leq R^{1}(1)-R^{1}(0).$
Then we want to prove the LHS of Eq. (C.24) under the same assumption.
$\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(1,s_{h+1}^{2}\right)\right)-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(0,s_{h+1}^{2}\right)\right)$
$=\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(1)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(0)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$=R^{1}(1)-R^{1}(0)+\left(q^{1}_{11}(1)-q^{1}_{01}(1)\right)\left(\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(f_{s_{h+1}^{2},h+1}(1)-f_{s_{h+1}^{2},h+1}(0)\right)\right)$
$\geq R^{1}(1)-R^{1}(0)+\left(q^{1}_{11}(1)-q^{1}_{01}(1)\right)\left(R^{1}(1)-R^{1}(0)\right)$
$\geq 0.$
This shows the LHS of Eq. (C.24). Using the same logic used in Case 1 will give us Eq. (C.25) for $t=h$. ∎
C.4 Proof of Theorem 3.6
See 3.6
Proof.
By Lemma 3.4, we only need to care which $B$ agents do we need to give action to. WLOG, assume that $I(1)\leq I(0)$. For simplicity, we first look at the simple case where $M=2,B=1$. The generalization to multi-agent case is straightforward and will be deferred to the last part of the proof. First of all, we will derive some useful equations that are useful in the proof.
$\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{0}_{1s_{h+1}^{1}}q^{1}_{0s_{h+1}^{2}}V_{h+1}^{*}\left(\left(s_{h+1}^{1},s_{h+1}^{2}\right)\right)-\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{1}_{1s_{h+1}^{1}}q^{0}_{0s_{h+1}^{2}}V_{h+1}^{*}\left(\left(s_{h+1}^{1},s_{h+1}^{2}\right)\right)$
$=q^{0}_{11}q^{1}_{01}V_{h+1}^{*}\left((1,1)\right)+q^{0}_{11}q^{1}_{00}V_{h+1}^{*}\left((1,0)\right)+q^{0}_{10}q^{1}_{01}V_{h+1}^{*}\left((1,0)\right)+q^{0}_{10}q^{1}_{00}V_{h+1}^{*}\left((0,0)\right)$
$-q^{1}_{11}q^{0}_{01}V_{h+1}^{*}\left((1,1)\right)-q^{1}_{11}q^{0}_{00}V_{h+1}^{*}\left((1,0)\right)-q^{1}_{10}q^{0}_{01}V_{h+1}^{*}\left((1,0)\right)-q^{1}_{10}q^{0}_{00}V_{h+1}^{*}\left((0,0)\right)$
$=q^{0}_{11}q^{1}_{01}V_{h+1}^{*}\left((1,1)\right)-q^{0}_{11}q^{1}_{01}V_{h+1}^{*}\left((1,0)\right)+q^{0}_{11}V_{h+1}^{*}\left((1,0)\right)$
$+q^{1}_{01}V_{h+1}^{*}\left((1,0)\right)-q^{0}_{11}q^{1}_{01}V_{h+1}^{*}\left((1,0)\right)$
$+V_{h+1}^{*}\left((0,0)\right)+q^{0}_{11}q^{1}_{01}V_{h+1}^{*}\left((0,0)\right)-\left(q^{0}_{11}+q^{1}_{01}\right)V_{h+1}^{*}\left((0,0)\right)$
$-q^{1}_{11}q^{0}_{01}V_{h+1}^{*}\left((1,1)\right)+q^{1}_{11}q^{0}_{01}V_{h+1}^{*}\left((1,0)\right)-q^{1}_{11}V_{h+1}^{*}\left((1,0)\right)$
$-q^{0}_{01}V_{h+1}^{*}\left((1,0)\right)+q^{1}_{11}q^{0}_{01}V_{h+1}^{*}\left((1,0)\right)$
$-V_{h+1}^{*}\left((0,0)\right)-q^{1}_{11}q^{0}_{01}V_{h+1}^{*}\left((0,0)\right)+\left(q^{1}_{11}+q^{0}_{01}\right)V_{h+1}^{*}\left((0,0)\right)$
$=\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right).$ (C.26)
Next, we will simplify the difference between the value function of different states, assuming that the optimal policy is the greedy policy. For simplicity, we will only give the detailed derivation of the equation for $V_{h}^{*}\left((1,1)\right)-V_{h}^{*}\left((1,0)\right)$. The equation for $V_{h}^{*}\left((1,0)\right)-V_{h}^{*}\left((0,0)\right)$ holds for similar reasons.
$V_{h}^{*}\left((1,1)\right)-V_{h}^{*}\left((1,0)\right)$ $=2R(1)+\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{1}_{1s_{h+1}^{1}}q\left(s_{h+1}^{2}\mid 1,0\right)V_{h+1}^{*}\left(\left(s_{h+1}^{1},s_{h+1}^{2}\right)\right)$
$-R(1)-R(0)-\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{0}_{1s_{h+1}^{1}}q^{1}_{0s_{h+1}^{2}}V_{h+1}^{*}\left(\left(s_{h+1}^{1},s_{h+1}^{2}\right)\right)$
$=R(1)-R(0)+q^{1}_{11}q^{0}_{11}V_{h+1}^{*}\left((1,1)\right)+q^{1}_{11}q^{0}_{10}V_{h+1}^{*}\left((1,0)\right)$
$+q^{1}_{10}q^{0}_{11}V_{h+1}^{*}\left((1,0)\right)+q^{0}_{10}q^{1}_{10}V_{h+1}^{*}\left((0,0)\right)$
$-q^{0}_{11}q^{1}_{01}V_{h+1}^{*}\left((1,1)\right)-q^{0}_{11}q^{1}_{00}V_{h+1}^{*}\left((1,0)\right)$
$-q^{0}_{10}q^{1}_{01}V_{h+1}^{*}\left((1,0)\right)-q^{0}_{10}q^{1}_{00}V_{h+1}^{*}\left((0,0)\right)$
$=R(1)-R(0)+q^{1}_{11}q^{0}_{11}V_{h+1}^{*}\left((1,1)\right)+q^{1}_{11}V_{h+1}^{*}\left((1,0)\right)-q^{1}_{11}q^{0}_{11}V_{h+1}^{*}\left((1,0)\right)$
$+q^{0}_{11}V_{h+1}^{*}\left((1,0)\right)-q^{1}_{11}q^{0}_{11}V_{h+1}^{*}\left((1,0)\right)$
$+V_{h+1}^{*}\left((0,0)\right)+q^{0}_{11}q^{1}_{11}V_{h+1}^{*}\left((0,0)\right)-\left(q^{1}_{11}+q^{0}_{11}\right)V_{h+1}^{*}\left((0,0)\right)$
$-q^{0}_{11}q^{1}_{01}V_{h+1}^{*}\left((1,1)\right)-q^{0}_{11}V_{h+1}^{*}\left((1,0)\right)+q^{1}_{01}q^{0}_{11}V_{h+1}^{*}\left((1,0)\right)$
$-q^{1}_{01}V_{h+1}^{*}\left((1,0)\right)+q^{0}_{11}q^{1}_{01}V_{h+1}^{*}\left((1,0)\right)$
$-V_{h+1}^{*}\left((0,0)\right)-q^{0}_{11}q^{1}_{01}V_{h+1}^{*}\left((0,0)\right)+\left(q^{0}_{11}+q^{1}_{01}\right)V_{h+1}^{*}\left((0,0)\right)$
$=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right).$ (C.27)
Similarly, we can have
$V_{h}^{*}\left((1,0)\right)-V_{h}^{*}\left((0,0)\right)$ $=R(1)-R(0)+\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right).$ (C.28)
We will derive the final useful equation for the incremental reward.
$I(1)$ $=\phi(1,1)-\phi(1,0)$
$=q^{1}_{11}R(1)+q^{1}_{10}R(0)-q^{0}_{11}R(1)-q^{0}_{10}R(0)$ (C.29)
$=\left(q^{1}_{11}-q^{0}_{11}\right)\left(R(1)-R(0)\right).$ (C.30)
Similarly, we also have
$I(0)=\left(q^{1}_{01}-q^{0}_{01}\right)\left(R(1)-R(0)\right).$ (C.31)
In the following, we will consider three cases.
Case 1:
$q^{0}_{11}\geq q^{0}_{01},q^{1}_{11}\leq q^{1}_{01}$. We will use backward induction to prove the following statements:
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (C.32)
$0\leq V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq R(1)-R(0)$ (C.33)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.34)
The statements hold trivially when $t=T$. Suppose the statements hold when $t\geq h+1$, when $t=h$, we will first prove Eq. (C.34) for $t=h$. Recall that the policy satisfying the following Bellman optimality equation is the optimal policy:
$V_{h}^{*}\left((s_{h}^{1},s_{h}^{2})\right)=R(s_{h}^{1})+R(s_{h}^{2})+\max_{a_{h+1}^{1},a_{h+1}^{2}}\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{a_{h+1}^{1}}_{s_{h}^{1}s_{h+1}^{1}}q^{a_{h+1}^{2}}_{s_{h}^{2}s_{h+1}^{2}}V_{h+1}^{*}\left(s_{h+1}^{1},s_{h+1}^{2}\right).$ (C.35)
By Lemma 3.4, the maximizer is attained when one of the $a_{h+1}^{i}=1$. Therefore, it’s easy to see that when $s_{h}^{1}=s_{h}^{2}$, randomly giving action will satisfy this equation. We will focus on the case where $s_{h}^{1}\neq s_{h}^{2}$. In that case, to show Eq. (C.34), we only need to show that
$\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{0}_{1s_{h+1}^{1}}q^{1}_{0s_{h+1}^{2}}V_{h+1}^{*}\left(s_{h+1}^{1},s_{h+1}^{2}\right)\geq\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{1}_{1s_{h+1}^{1}}q^{0}_{0s_{h+1}^{2}}V_{h+1}^{*}\left(s_{h+1}^{1},s_{h+1}^{2}\right).$ (C.36)
By Eq. (C.26), this is equivalent to showing
$\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)\geq 0.$
Then we have
$\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$-\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$=\frac{I(0)-I(1)}{R(1)-R(0)}\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)-\frac{\left(q^{1}_{11}I(0)-q^{1}_{01}I(1)\right)}{R(1)-R(0)}\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$=\frac{\left(q^{1}_{10}I(0)-q^{1}_{00}I(1)\right)}{R(1)-R(0)}\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq 0,$
where the first inequality holds because of the induction hypothesis and $q^{0}_{11}\geq q^{0}_{01},q^{1}_{01}\geq q^{1}_{11}$, the first equality holds because of Eq. (C.31) and Eq. (C.30) and the last inequality holds because $q^{1}_{10}\geq q^{1}_{00}$, $I(0)\geq I(1)$ and the assumption that Eq. (C.32) holds for $t=h+1$.
Next, we first show the LHS of Eq. (C.32) for $t=h$ under the assumption that Eq. (C.34) holds for $t=h$ and Eq. (C.32) holds for $t=h+1$.
$V_{h}^{*}\left((1,0)\right)-V_{h}^{*}\left((0,0)\right)$ $=R(1)-R(0)+\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq R(1)-R(0)+\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$-q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq R(1)-R(0)+q^{1}_{00}\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq 0,$
where the first equality holds because of Eq. (C.28) and the last inequality holds since $q^{0}_{11}\geq q^{0}_{01}$. Under the same condition, for the RHS of Eq. (C.32), we have
$q^{0}_{10}\left(V_{h}^{*}\left((1,0)\right)-V_{h}^{*}\left((0,0)\right)\right)$ $=q^{0}_{10}(R(1)-R(0))$
$+q^{0}_{10}q^{1}_{00}\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{10}q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-V_{h+1}^{*}\left((1,0)\right)\right)$
$\leq q^{0}_{10}(R(1)-R(0))$
$+q^{1}_{00}\left(q^{0}_{11}-q^{0}_{01}\right)(R(1)-R(0))$
$+q^{0}_{10}q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)(R(1)-R(0))$
$\leq q^{0}_{10}(R(1)-R(0))$
$+q^{1}_{00}q^{0}_{11}(R(1)-R(0))+q^{1}_{01}q^{0}_{11}(R(1)-R(0))$
$=R(1)-R(0).$
Then, we only need to show Eq. (C.33) for $t=h$. Assuming that Eq. (C.34) holds for $t=h$ and Eq. (C.33) holds for $t=h+1$, for the RHS, we have
$V_{h}^{*}\left((1,1)\right)-V_{h}^{*}\left((1,0)\right)$ $=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-V_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{0}_{10}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$\leq R(1)-R(0),$
where the last inequality holds since $q^{1}_{11}\leq q^{1}_{01}$. Under the same condition, for the LHS, we have
$V_{h}^{*}\left((1,1)\right)-V_{h}^{*}\left((1,0)\right)$ $=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-V_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{0}_{10}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq R(1)-R(0)+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)(R(1)-R(0))$
$+\left(q^{1}_{11}-q^{1}_{01}\right)(R(1)-R(0))$
$=\left(q^{1}_{00}+q^{1}_{11}-q^{0}_{11}q^{1}_{01}+q^{0}_{11}q^{1}_{01}\right)(R(1)-R(0))$
$\geq 0,$
where the last inequality holds since $q^{1}_{11}\geq q^{0}_{11}$ by Lemma 3.2.
Case 2:
$q^{0}_{11}\leq q^{0}_{01},q^{1}_{11}\leq q^{1}_{01}$.
First we consider the sub-case where $\boxed{\bm{q^{1}_{11}q^{0}_{01}\leq q^{1}_{01}q^{0}_{11}}}$. We will use backward induction to prove the following statements:
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))\leq R(1)-R(0)$ (C.37)
$0\leq V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq R(1)-R(0)$ (C.38)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.39)
When $t=T$, the statements hold trivially. Suppose the statements hold for $t\geq h+1$, when $t=h$, we will first prove Eq. (C.39) under the assumption that Eq. (C.37) and Eq. (C.38) holds for $t=h+1$. Similarly we only need to show Eq. (C.36).
$\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq\frac{\left(q^{1}_{10}I(0)-q^{1}_{00}I(1)\right)}{R(1)-R(0)}\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq 0,$
where the first inequality holds because the assumption we make and the last inequality holds because $q^{1}_{10}\geq q^{1}_{00}$, $I(0)\geq I(1)$ and the assumption that Eq. (C.37) holds.
Next we want to prove the RHS of Eq. (C.38) under the assumption that Eq. (C.39) holds for $t=h$ and that Eq. (C.37) and Eq. (C.38) hold for $t=h+1$
$V_{h}^{*}\left((1,1)\right)-V_{h}^{*}\left((1,0)\right)$ $=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-V_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{0}_{10}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$\leq R(1)-R(0).$
Under the same condition, for the LHS of Eq. (C.38), we have
$V_{h}^{*}\left((1,1)\right)-V_{h}^{*}\left((1,0)\right)$ $=R(1)-R(0)+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-V_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{0}_{10}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq R(1)-R(0)+q^{0}_{10}\left(q^{1}_{11}-q^{1}_{01}\right)(R(1)-R(0))$
$+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)(R(1)-R(0))$
$=(q^{1}_{11}+q^{1}_{00})(R(1)-R(0))$
$\geq 0.$
Similarly, we can show Eq. (C.37) for $t=h$ under the same condition.
Secondly, we consider the sub-case where $\boxed{\bm{q^{1}_{11}q^{0}_{01}\geq q^{1}_{01}q^{0}_{11}}}$. We want to use backward induction to prove the following statements:
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))\leq{R(1)-R(0)}$ (C.40)
$0\leq V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq R(1)-R(0)$ (C.41)
$V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (C.42)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.43)
When $t=T$, the statements hold trivially. Suppose this holds for $t\geq h+1$, when $t=h$, we will first show Eq. (C.43) for $t=h$ under the assumption that Eq. (C.40), Eq. (C.41) and Eq. (C.42) hold for $t=h+1$. Similarly, this is equivalent to show Eq. (C.36).
$\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$=\frac{I(0)-I(1)}{R(1)-R(0)}\left(V_{h+1}^{*}(1,0)-V_{h+1}^{*}(0,0)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq 0,$
where the first equality holds because of Eq. (C.31) and Eq. (C.30) and the first inequality holds since $q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\leq 0$ and the assumption that Eq. (C.42) holds for $t=h+1$.
Using the same logic as the previous case, we can prove that
$0\leq V_{h}^{*}\left((1,0)\right)-V_{h}^{*}((0,0))\leq{R(1)-R(0)}$
$0\leq V_{h}^{*}\left((1,1)\right)-V_{h}^{*}((1,0))\leq R(1)-R(0).$
Therefore, the only thing we need to show is the Eq. (C.42) for $t=h$ under the assumption that Eq. (C.43) holds for $t=h$ and that Eq. (C.42) holds for $t=h+1$. We have
$V_{h}^{*}\left((1,1)\right)-2V_{h}^{*}((1,0))+V_{h}^{*}((0,0))$
$=q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}((1,0))+V_{h+1}^{*}((0,0))\right)$
$+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$-q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}((1,0))+V_{h+1}^{*}((0,0))\right)$
$-\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$=q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}((1,0))+V_{h+1}^{*}((0,0))\right)$
$-q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}((1,0))+V_{h+1}^{*}((0,0))\right)$
$+\frac{I(1)-I(0)}{R(1)-R(0)}\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right).$
Since $I(1)\leq I(0),V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\geq 0$, we only need to prove that
$q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\geq q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right).$
Since we have
$q^{1}_{11}q^{0}_{01}\geq q^{1}_{01}q^{0}_{11},$
we have
$q^{1}_{11}q^{0}_{01}-q^{1}_{11}q^{0}_{11}$ $\geq q^{1}_{01}q^{0}_{11}-q^{1}_{11}q^{0}_{11}$
$q^{1}_{11}\left(q^{0}_{01}-q^{0}_{11}\right)$ $\geq q^{0}_{11}(q^{1}_{01}-q^{1}_{11})$
$q^{1}_{01}\left(q^{0}_{01}-q^{0}_{11}\right)\geq q^{1}_{11}\left(q^{0}_{01}-q^{0}_{11}\right)$ $\geq q^{0}_{11}(q^{1}_{01}-q^{1}_{11})$
$q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)$ $\leq q^{0}_{11}(q^{1}_{11}-q^{1}_{01}).$
This will give us the desired result.
Case 3:
$q^{1}_{11}\geq q^{1}_{01},q^{0}_{11}\geq q^{0}_{01}$.
We will use backward induction to prove the following statements:
$\forall t\leq T,$ $\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (C.44)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (C.45)
$\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (C.46)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (C.47)
$\pi_{t}^{\text{greedy}}=\pi_{t}^{*}.$ (C.48)
When $t=T$, this holds trivially, suppose this holds for $t\geq h+1$. When $t=h$, we will first show that Eq. (C.48) holds for $t=h$ under the assumption that Eq. (C.44), Eq. (C.45), Eq. (C.46) and Eq. (C.47) hold for $t=h+1$. This again is equivalent to show that
$\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)\geq 0.$
We have
$\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$=\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}-q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-V_{h+1}^{*}\left((1,0)\right)\right).$
If $q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}-q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\geq 0$, then using our assumption that, $V_{h+1}^{*}\left((1,1)\right)-V_{h+1}^{*}\left((1,0)\right)\geq 0,V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\geq 0$, will give us the desired result.
If $q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}-q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\leq 0$, with our hypothesis, we have
$\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}-q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-V_{h+1}^{*}\left((1,0)\right)\right)$
$\geq\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}-q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)$
×$\times\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-h}\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h}\right).$ (C.49)
Denote $\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-h}\right)$ as $d_{\text{upper}}$ and let $k$ be defined as
$\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h}\right)/d_{\text{upper}}$. Then to show that Eq. (C.49) is larger than 0, we only need to show that
$q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}-(k-1)\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\geq 0,$
which is equivalent to show $k\geq\frac{q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}-q^{0}_{11}-q^{1}_{01}+q^{1}_{11}+q^{0}_{01}}{q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}}$.
First, we will show
$\frac{q^{0}_{10}+q^{0}_{01}}{q^{1}_{10}+q^{1}_{01}}\geq\frac{q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}-q^{0}_{11}-q^{1}_{01}+q^{1}_{11}+q^{0}_{01}}{q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}}.$
To do this, we only need to show
$\left(q^{0}_{10}+q^{0}_{01}\right)\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)$
$-\left(q^{1}_{10}+q^{1}_{01}\right)\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)$
$-\left(q^{1}_{10}+q^{1}_{01}\right)\left(q^{1}_{11}+q^{0}_{01}-q^{0}_{11}-q^{1}_{01}\right)\geq 0.$
We have
$\left(q^{0}_{10}+q^{0}_{01}\right)\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)$
$-\left(q^{1}_{10}+q^{1}_{01}\right)\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)$
$-\left(q^{1}_{10}+q^{1}_{01}\right)\left(q^{1}_{11}+q^{0}_{01}-q^{0}_{11}-q^{1}_{01}\right)$
$=\left(q^{0}_{10}+q^{0}_{01}-q^{1}_{10}-q^{1}_{01}\right)\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)$
$-\left(q^{1}_{10}+q^{1}_{01}\right)\left(q^{1}_{11}+q^{0}_{01}-q^{0}_{11}-q^{1}_{01}\right)$
$=\left(q^{1}_{11}+q^{0}_{01}-q^{0}_{11}-q^{1}_{01}\right)\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}-q^{1}_{10}-q^{1}_{01}\right).$
Since $q^{1}_{11}+q^{0}_{01}-q^{0}_{11}-q^{1}_{01}=\frac{I(1)-I(0)}{R(1)-R(0)}\leq 0$ and $q^{0}_{11}q^{1}_{01}-q^{1}_{01}\leq 0$, we have the desired result.
Second, we want to show that $1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h}\geq 1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-h}$. To do this, we only need to show that $\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h}\leq\left(q^{0}_{11}-q^{0}_{01}\right)^{T-h}$, which comes directly from the fact that $0\leq q^{1}_{11}-q^{1}_{01}\leq q^{0}_{11}-q^{0}_{01}$.
Then we have
$k$ $=\frac{q^{0}_{10}+q^{0}_{01}}{q^{1}_{10}+q^{1}_{01}}\frac{1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h}}{1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-h}}$
$\geq\frac{q^{0}_{10}+q^{0}_{01}}{q^{1}_{10}+q^{1}_{01}}$
$\geq\frac{q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}-q^{0}_{11}-q^{1}_{01}+q^{1}_{11}+q^{0}_{01}}{q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}}.$
This will give us the desired result.
Next, we will first prove Eq. (C.44) for $t=h$ under the assumption that Eq. (C.44), Eq. (C.45), Eq. (C.46) and Eq. (C.47) hold for $t=h+1$ and Eq. (C.43) holds for $t=h$ We have
$V_{h}^{*}\left((1,1)\right)-V_{h}^{*}\left((1,0)\right)$ $=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-V_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{0}_{10}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$\geq R(1)-R(0)$
$+\left(q^{1}_{11}-q^{1}_{01}\right)\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h}\right)$
$=\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h+1}\right).$
Under the same condition, for Eq. (C.45) when $t=h$, we have
$V_{h}^{*}\left((1,1)\right)-V_{h}^{*}\left((1,0)\right)$ $=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-V_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{0}_{10}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,0)\right)-V_{h+1}^{*}\left((0,0)\right)\right)$
$\leq R(1)-R(0)$
$+\left(q^{0}_{11}-q^{0}_{01}\right)\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-h}\right)$
$=\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-h+1}\right),$
where the inequality holds because we have $q^{1}_{11}-q^{1}_{01}\leq q^{0}_{11}-q^{0}_{01}$. Using the same logic will give us the Eq. (C.47) and Eq. (C.46) for $t=h$.
Now we are going to prove the theorem for arbitrary number of agents and budget. We will again prove the result under three cases.
Case 1:
Suppose $q^{1}_{11}\leq q^{1}_{01},q^{0}_{11}\geq q^{0}_{01}$.
We will first define three sets: $\mathcal{S}^{M-2}_{1},\mathcal{S}^{M-2}_{2},\mathcal{S}^{M-2}_{3}$.
$\mathcal{S}^{M-2}_{1}$ :=$:=\{\bm{s}:\bm{s}\in\mathcal{S}^{M-2},n_{\bm{s}}\geq B\}$
$\mathcal{S}^{M-2}_{2}$ :=$:=\{\bm{s}:\bm{s}\in\mathcal{S}^{M-2},n_{\bm{s}}=B-1\}$
$\mathcal{S}^{M-2}_{3}$ :=$:=\{\bm{s}:\bm{s}\in\mathcal{S}^{M-2},n_{\bm{s}}\leq B-2\},$
where $n_{\bm{s}}$ denotes the number of state 00 of $\bm{s}$.
We will use backward induction to prove that
$\forall t\leq T,\forall\bm{s}\in\mathcal{S}^{M-2}_{3},$ $0\leq V_{t}^{*}((1,0,\bm{s}))-V_{t}^{*}((0,0,\bm{s}))\leq R(1)-R(0)$ (C.50)
$0\leq V_{t}^{*}((1,1,\bm{s}))-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (C.51)
$\forall\bm{s}\in\mathcal{S}^{M-2}_{2},$ $0\leq V_{t}^{*}((1,0,\bm{s}))-V_{t}^{*}((0,0,\bm{s}))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (C.52)
$0\leq V_{t}^{*}((1,1,\bm{s}))-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (C.53)
$\forall\bm{s}\in\mathcal{S}^{M-2}_{1},$ $0\leq V_{t}^{*}((1,0,\bm{s}))-V_{t}^{*}((0,0,\bm{s}))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (C.54)
$0\leq V_{t}^{*}((1,1,\bm{s}))-V_{t}^{*}((1,0,\bm{s}))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (C.55)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.56)
This holds trivially when $t=T$. Suppose $t\geq h+1$, this holds, we will first prove Eq. (C.56) for $t=h$ under the assumption that Eq. (C.50), Eq. (C.51), Eq. (C.52), Eq. (C.53), Eq. (C.54) and Eq. (C.55) hold for $t=h+1$. Similarly we have the following Bellman optimality equation.
$V_{h}^{*}\left(\left(1,0,\bm{s}_{h}\right)\right)$ $=R(1)+R(0)+\sum_{s\in\bm{s}_{h}}R(s)$
$+\max_{a_{h+1}^{1},a_{h+1}^{2},\cdots,a_{h+1}^{M}}\sum_{s_{h+1}^{1},s_{h+1}^{2}}q_{1,s_{h+1}^{1}}^{a_{h+1}^{1}}q_{0,s_{h+1}^{2}}^{a_{h+1}^{2}}\sum_{\bm{s}_{h+1}}q_{\bm{s}_{h},\bm{s}_{h+1}}^{\bm{a}_{h+1}}V_{h+1}^{*}\left((s_{h+1}^{1},s_{h+1}^{2},\bm{s}_{h+1})\right).$
We want to show that fix any policy of choosing $a_{h+1}^{3},\cdots,a_{h+1}^{M}$, giving action 1 to agent who is in state 0 will give a higher value, which is
$\forall\bm{a}_{h+1},$ $\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{0}_{1s_{h+1}^{1}}q^{1}_{0s_{h+1}^{2}}\sum_{\bm{s}_{h+1}}q_{\bm{s}_{h},\bm{s}_{h+1}}^{\bm{a}_{h+1}}V_{h+1}^{*}\left((s_{h+1}^{1},s_{h+1}^{2},\bm{s}_{h+1})\right)$
$\geq\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{1}_{1s_{h+1}^{1}}q^{0}_{0s_{h+1}^{2}}\sum_{\bm{s}_{h+1}}q_{\bm{s}_{h},\bm{s}_{h+1}}^{\bm{a}_{h+1}}V_{h+1}^{*}\left((s_{h+1}^{1},s_{h+1}^{2},\bm{s}_{h+1})\right).$ (C.57)
Suppose Eq. (C.57) holds, starting from any policy of allocating action that is not greedy, we can find an agent who is in state 1 and gets action 1 while another agent who is in state 0 gets action 0. By using the result, switching the action between them will give us a better policy. Suppose there are less than $B$ agents who are in state 0, then we can keep switching until all agents who are in state 0 are given action 1 , this is the greedy policy. Suppose there are more than $B$ agents who are in state 0, then we can keep switching until $B$ agents who are in state 0 get action 1, which is the greedy policy.
Now, we need to prove Eq. (C.57). Eq. (C.57) is equivalent to
$\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(F_{\bm{a}_{h+1},h+1}((1,1))-F_{\bm{a}_{h+1},h+1}((1,0))\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,1)\right)-2F_{\bm{a}_{h+1},h+1}\left((1,0)\right)+F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)\geq 0,$ (C.58)
where $F_{\bm{a}_{h+1},h+1}\left((s_{h+1}^{1},s_{h+1}^{2})\right)=\sum_{\bm{s}_{h+1}}q_{\bm{s}_{h},\bm{s}_{h+1}}^{\bm{a}_{h+1}}V_{h+1}^{*}\left((s_{h+1}^{1},s_{h+1}^{2},\bm{s}_{h+1})\right)$. The reason why this holds is similar to what we proved in binary agent case. We know from the assumption that Eq. (C.50)-(C.55) hold and the definition of $F$ that
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,1))-F_{\bm{a}_{h+1},h+1}((1,0))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,0))-F_{\bm{a}_{h+1},h+1}((0,0))\leq\frac{R(1)-R(0)}{q^{0}_{10}}.$
Then, we have
$\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,1)\right)-2F_{\bm{a}_{h+1},h+1}\left((1,0)\right)+F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$\geq\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$-\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$=\frac{I(0)-I(1)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)-\frac{\left(q^{1}_{11}I(0)-q^{1}_{01}I(1)\right)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$=\frac{\left(q^{1}_{10}I(0)-q^{1}_{00}I(1)\right)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$\geq 0.$
This shows Eq. (C.57), which, by our argument, shows Eq. (C.56) for $t=h$. Next, we will first define the specific greedy policy instead of randomly choosing agents who have higher incremental reward at time $h$ for different $\bm{s}$. When $\bm{s}\in\mathcal{S}^{M-2}_{1}$, we will only give action 1 to agent $3,4,\cdots,M$; when $\bm{s}\in\mathcal{S}^{M-2}_{2}$, we will give one action 1 to agent $1,2$; when $\bm{s}\in\mathcal{S}^{M-2}_{3}$, we will give two action 1 to agent $1,2$.
For $\bm{s}\in\mathcal{S}_{1}^{M-2}$, we will show Eq. (C.54) holds for $t=h$ under the assumption that Eq.(C.50)-(C.55) hold for $t=h+1$ and Eq. (C.56) holds for $t=h$. We have
$V_{h}^{*}\left((1,0,\bm{s})\right)-V_{h}^{*}\left((0,0,\bm{s})\right)$ $=R(1)-R(0)+\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{0}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-F_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{0}_{00}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$\geq 0.$
Next, we will show the RHS of Eq. (C.54) for $t=h$ under the same condition.
$q_{10}^{0}\left(V_{h}^{*}\left((1,0,\bm{s})\right)-V_{h}^{*}\left((0,0,\bm{s})\right)\right)$ $=q_{10}^{0}(R(1)-R(0))+q_{10}^{0}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$+q_{10}^{0}q^{0}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$=q_{10}^{0}(R(1)-R(0))+q^{0}_{10}q^{0}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-F_{h+1}^{*}\left((1,0)\right)\right)$
$+q_{10}^{0}q^{0}_{00}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$\leq q_{10}^{0}(R(1)-R(0))+q_{01}^{0}q_{11}^{0}(R(1)-R(0))+q_{00}^{0}q_{11}^{0}(R(1)-R(0))$
$=R(1)-R(0).$
Meanwhile, when $t=h$, again under the same condition, the LHS of Eq. (C.55) satisfies
$V_{h}^{*}\left((1,1,\bm{s})\right)-V_{h}^{*}\left((1,0,\bm{s})\right)$ $=R(1)-R(0)+\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{11}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{0}_{11}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-F_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{0}_{10}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$\geq 0.$
With the same reasoning, we can show the RHS of Eq. (C.55) under the same condition for $t=h$.
For $\bm{s}\in\mathcal{S}_{3}^{M-2}$, we have
$V_{h}^{*}\left((1,0,\bm{s})\right)-V_{h}^{*}\left((0,0,\bm{s})\right)$ $=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{1}_{01}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{1}_{01}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-F_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{1}_{00}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$=\frac{q^{0}_{10}}{q^{0}_{10}}\left(R(1)-R(0)+q^{1}_{01}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-F_{h+1}^{*}\left((1,0)\right)\right)\right)$
$+\frac{q^{0}_{10}}{q^{0}_{10}}\left(q^{1}_{00}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)\right)$
$\geq\frac{q^{0}_{10}}{q^{0}_{10}}\left(R(1)-R(0)+\frac{q^{1}_{01}}{q^{0}_{10}}\left(q^{1}_{11}-q^{1}_{01}\right)\left(R(1)-R(0)\right)\right)$
$+\frac{q^{1}_{00}}{q^{0}_{10}}\left(q^{1}_{11}-q^{1}_{01}\right)\left(R(1)-R(0)\right)$
$=\frac{1}{q^{0}_{10}}\left(q^{0}_{10}+q^{1}_{00}q^{1}_{11}-q^{1}_{00}q^{1}_{01}\right)\left(R(1)-R(0)\right)$
$+\frac{1}{q^{0}_{10}}\left(q^{1}_{01}\left(q^{1}_{11}-q^{1}_{01}\right)\right)\left(R(1)-R(0)\right)$
$\geq\frac{1}{q^{0}_{10}}\left(1-q^{1}_{01}+q^{1}_{11}-q^{0}_{11}\right)\left(R(1)-R(0)\right)$
$\geq 0.$
Meanwhile, we also have
$V_{h}^{*}\left((1,1,\bm{s})\right)-V_{h}^{*}\left((1,0,\bm{s})\right)$ $=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{1}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{1}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-F_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{1}_{10}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$=\frac{q^{0}_{10}}{q^{0}_{10}}\left(R(1)-R(0)+q^{1}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-F_{h+1}^{*}\left((1,0)\right)\right)\right)$
$+\frac{q^{0}_{10}}{q^{0}_{10}}\left(q^{1}_{10}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)\right)$
$\geq\frac{q^{0}_{10}}{q^{0}_{10}}\left(R(1)-R(0)+\frac{q^{1}_{11}}{q^{0}_{10}}\left(q^{1}_{11}-q^{1}_{01}\right)\left(R(1)-R(0)\right)\right)$
$+\frac{q^{1}_{10}}{q^{0}_{10}}\left(q^{1}_{11}-q^{1}_{01}\right)\left(R(1)-R(0)\right)$
$=\frac{1}{q^{0}_{10}}\left(q^{0}_{10}+q^{1}_{10}q^{1}_{11}-q^{1}_{10}q^{1}_{01}\right)\left(R(1)-R(0)\right)$
$+\frac{1}{q^{0}_{10}}\left(q^{1}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\right)\left(R(1)-R(0)\right)$
$\geq\frac{1}{q^{0}_{10}}\left(1-q^{1}_{01}+q^{1}_{11}-q^{0}_{11}\right)\left(R(1)-R(0)\right)$
$\geq 0.$
For $\bm{s}\in\mathcal{S}_{2}^{M-2}$, we can use the same reasoning used in the two agent case to get the result.
Case 2:
$q^{0}_{11}\leq q^{0}_{01},q^{1}_{11}\leq q^{1}_{01}$.
First we consider the sub-case where $\boxed{\bm{q^{1}_{11}q^{0}_{01}\leq q^{1}_{01}q^{0}_{11}}}$. We will use backward induction to prove the following statements:
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))\leq R(1)-R(0)$ (C.59)
$0\leq V_{t}^{*}\left((1,1,\bm{s})\right)-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (C.60)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.61)
They hold trivially when $t=T$. Suppose they hold when $t\geq h+1$, when $t=h$, we will first prove Eq. (C.61) for $t=h$ under the assumption that Eq. (C.59) and Eq. (C.38) hold for $t=h+1$. Using the same reasoning, we only need to show Eq. (C.58). Again, from the assumptions and the definition of $F$, we know that
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,1))-F_{\bm{a}_{h+1},h+1}((1,0))\leq{R(1)-R(0)}$
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,0))-F_{\bm{a}_{h+1},h+1}((0,0))\leq{R(1)-R(0)}.$
Therefore, we have
$\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,1)\right)-2F_{\bm{a}_{h+1},h+1}\left((1,0)\right)+F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$\geq\left(q^{0}_{11}+q^{1}_{01}-q^{1}_{11}-q^{0}_{01}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$-\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$=\frac{I(0)-I(1)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)-\frac{\left(q^{1}_{11}I(0)-q^{1}_{01}I(1)\right)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$=\frac{\left(q^{1}_{10}I(0)-q^{1}_{00}I(1)\right)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
$\geq 0.$
This will give us the desired result. Assuming Eq. (C.61) holds for $t=h$, we will select the greedy policy for $\mathcal{S}_{i}^{M-2}$ as we did in case 1. Again, the only case that is different from the 2-agent case is when $\bm{s}\in\mathcal{S}_{1}^{M-2},\mathcal{S}_{3}^{M-2}$. Therefore, we will only show Eq. (C.59) and Eq. (C.60) for those two cases when $t=h$ under the assumption that Eq. (C.59) and Eq. (C.60) hold for $t=h+1$ and Eq. (C.61) holds for $t=h$. We will first prove the RHS of Eq. (C.59). When $\bm{s}\in\mathcal{S}_{1}^{M-2}$, we have
$V_{h}^{*}\left((1,0,\bm{s})\right)-V_{h}^{*}\left((0,0,\bm{s})\right)$ $=R(1)-R(0)+\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{0}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-F_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{0}_{00}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$\leq R(1)-R(0).$
Similarly, when $\bm{s}\in\mathcal{S}_{3}^{M-2}$, we have
$V_{h}^{*}\left((1,0,\bm{s})\right)-V_{h}^{*}\left((0,0,\bm{s})\right)$ $=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{1}_{01}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{1}_{01}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-F_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{1}_{00}\left(q^{1}_{11}-q^{1}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$\leq R(1)-R(0).$
Now, we are going to prove the LHS of Eq. (C.59). When $\bm{s}\in\mathcal{S}_{1}^{M-2}$, we have
$V_{h}^{*}\left((1,0,\bm{s})\right)-V_{h}^{*}\left((0,0,\bm{s})\right)$ $=R(1)-R(0)+\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$=R(1)-R(0)+q^{0}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-F_{h+1}^{*}\left((1,0)\right)\right)$
$+q^{0}_{00}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$\geq R(1)-R(0)+q^{0}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)(R(1)-R(0))+q^{0}_{00}\left(q^{0}_{11}-q^{0}_{01}\right)(R(1)-R(0))$
$=(1+q_{11}^{0}+q_{00}^{0})(R(1)-R(0))$
$\geq 0.$
Using the same calculation, we can show the case for $\bm{s}\in\mathcal{S}_{3}^{M-2}$ and Eq. (C.60).
Secondly, we consider the sub-case where $\boxed{\bm{q^{1}_{11}q^{0}_{01}\geq q^{1}_{01}q^{0}_{11}}}$. We want to use backward induction to prove the following statements:
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))\leq{R(1)-R(0)}$ (C.62)
$0\leq V_{t}^{*}\left((1,1,\bm{s})\right)-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (C.63)
$V_{t}^{*}\left((1,1,\bm{s})\right)-V_{t}^{*}((1,0,\bm{s}))\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (C.64)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.65)
When $t=T$, the statements hold trivially. Suppose this holds for $t\geq h+1$, when $t=h$, we will first show Eq. (C.65) under the assumption that Eq. (C.62)-(C.64) hold for $t=h+1$. Using the same reasoning, we only need to show Eq. (C.58). Again, from the above assumptions and the definition of $F$, we know that
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,1))-F_{\bm{a}_{h+1},h+1}((1,0))\leq{R(1)-R(0)}$
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,0))-F_{\bm{a}_{h+1},h+1}((0,0))\leq{R(1)-R(0)}.$
Then combined with the fact that ${{q^{1}_{11}q^{0}_{01}\geq q^{1}_{01}q^{0}_{11}}}$, it is easy to see that Eq. (C.58) holds.
Using the same logic used in the first sub-case, we can prove Eq. (C.62),(C.63) for $t=h$. The only thing we need to show is Eq. (C.64) for $t=h$. Again, when $\bm{s}\in\mathcal{S}_{2}^{M-2}$, this is the same as the two agent case. Therefore, we will only prove the other two cases under the assumption that Eq. (C.62)-(C.64) hold for $t=h+1$ and Eq. (C.65) holds for $t=h$. We will prove the result for $\bm{s}\in\mathcal{S}_{1}^{M-2}$, the other cases will hold for similar reasons. We have
$V_{h}^{*}\left((1,1,\bm{s})\right)-2V_{h}^{*}((1,0,\bm{s}))+V_{h}^{*}((0,0,\bm{s}))$
$=\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$+q^{0}_{11}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$-\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,0)\right)-F_{h+1}^{*}\left((0,0)\right)\right)$
$-q^{0}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$=\left(q^{0}_{11}-q^{0}_{01}\right)^{2}\left(F_{h+1}^{*}\left((1,1)\right)-2F_{h+1}^{*}\left((1,0)\right)+F_{h+1}^{*}\left((0,0)\right)\right)$
$\leq 0,$
where the last inequality holds because of the assumption that Eq. (C.64) holds for $t=h+1$.
Case 3:
$q^{1}_{11}\geq q^{1}_{01},q^{0}_{11}\geq q^{0}_{01}$. Similar to the two agent case, we can prove the following statements using backward induction:
$\forall t\leq T,$ $\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (C.66)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (C.67)
$\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (C.68)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (C.69)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.70)
The proof will be exactly using the idea used in 2 agent case and the calculation for multiple agents case discussed above, thus will be omitted. ∎
Appendix D Technical Details In Section 4
We will continue to use the notation defined in Appendix C.
D.1 Proof of Lemma 4.3
See 4.3
Proof.
By Bellman equation, we have
$V^{\bm{\pi}^{b}}_{0}\left(\bm{s}_{0}\right)=\mathbb{E}\left[\sum_{m=1}^{M}c_{0}^{m}\right]+\mathbb{E}_{\pi_{1}^{b}}\left[V^{\bm{\pi}^{b}}_{1}(\bm{s}_{1})\right].$
Since $V^{\bm{\pi}^{b}}_{0}\left(\bm{s}_{0}\right)=0,V^{\bm{\pi}^{b}}_{1}(\bm{s}_{1})\geq 0,P_{\bm{\pi}^{b}}(\bm{s}_{0},\bm{s}_{1})>0$, we have $V^{\bm{\pi}^{b}}_{1}(\bm{s}_{1})=0$ for all $\bm{s}_{1}\in\mathcal{S}^{M}$. Since again by Bellman equation, we have $V^{\bm{\pi}^{b}}_{1}(\bm{s}_{1})\geq\sum_{m=1}^{M}c_{1}^{m}\left(\pi_{2}^{b}(s_{1}^{m})\right)$, we have this sum should be 0. This will indicate that there are enough good arms for state $\bm{s}_{1}$. This will give us the desired result. ∎
Before proving Theorem 4.4, we first introduce Lemma D.1 which decomposes the cumulative regret to per-timestep regret.
Lemma D.1.
For any $t\leq T$, let per-timestep regret $\text{Regret}_{t}(\pi,\mathcal{H}_{t})$ for the policy $\bm{\pi}$ be defined as:
$\text{Regret}_{t}(\bm{\pi},\mathcal{H}_{t})$ $=\mathbb{E}_{\pi_{t+1}}\left[\sum_{m=1}^{M}c_{t}^{m}(a_{t+1}^{m})\mid\mathcal{H}_{t}\right]+\mathbb{E}_{\pi_{t+1}}\left[\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{k=t+1}^{T}c_{k}^{m}\left(a_{t+1}^{m}\right)|\bm{r}_{t+1},\bm{s}_{t+1},\bm{a}_{t+1},\mathcal{H}_{t}\right]\right]$
$-\mathbb{E}_{\pi_{t+1}^{b}}\left[\sum_{m=1}^{M}c_{t}^{m}\left(a_{t+1}^{m}\right)\mid\mathcal{H}_{t}\right]-\mathbb{E}_{\pi_{t+1}^{b}}\left[\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{k=t+1}^{T}c_{k}^{m}\left(a_{t+1}^{m}\right)|\bm{r}_{t+1},\bm{s}_{t+1},\bm{a}_{t+1},\mathcal{H}_{t}\right]\right].$
Then the cumulative regret $\mathcal{R}(\pi)$ can be decomposed as following:
$\mathcal{R}(\pi)=\mathbb{E}\left[\sum_{t=0}^{T}\text{Regret}_{t}\left(\pi,\mathcal{H}_{t}\right)\right].$
D.2 Proof of Lemma D.1
Proof.
We will prove the result iteratively from $t=1$ to $T$. All we need is calculation. We have
$\mathcal{R}(\bm{\pi})$ $=\mathbb{E}_{\bm{\pi}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}c_{t}^{m}(a_{t+1}^{m})|\bm{s}_{0}\right]-\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}c_{t}^{m}(a_{t+1}^{m})|\bm{s}_{0}\right]$
$=\mathbb{E}_{\pi_{1}}\left[\sum_{m=1}^{M}c_{0}^{m}(a_{1}^{m})\mid\mathcal{H}_{0}\right]+\mathbb{E}_{\pi_{1}}\left[\mathbb{E}_{\bm{\pi}}\left[\sum_{m=1}^{M}\sum_{k=1}^{T}c_{k}^{m}\left(a_{k+1}^{m}\right)|\bm{r}_{1},\bm{s}_{1},\bm{a}_{1},\mathcal{H}_{0}\right]\right]$
$-\mathbb{E}_{\pi_{1}}\left[\sum_{m=1}^{M}c_{0}^{m}(a_{1}^{m})\mid\mathcal{H}_{0}\right]-\mathbb{E}_{\pi_{1}}\left[\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{k=1}^{T}c_{k}^{m}\left(a_{k+1}^{m}\right)|\bm{r}_{1},\bm{s}_{1},\bm{a}_{1},\mathcal{H}_{0}\right]\right]$
$+\mathbb{E}_{\pi_{1}}\left[\sum_{m=1}^{M}c_{0}^{m}(a_{1}^{m})\mid\mathcal{H}_{0}\right]+\mathbb{E}_{\pi_{1}}\left[\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{k=1}^{T}c_{k}^{m}\left(a_{k+1}^{m}\right)|\bm{r}_{1},\bm{s}_{1},\bm{a}_{1},\mathcal{H}_{0}\right]\right]$
$-\mathbb{E}_{\pi_{1}^{b}}\left[\sum_{m=1}^{M}c_{0}^{m}\left(a_{1}^{m}\right)\mid\mathcal{H}_{0}\right]-\mathbb{E}_{\pi_{1}^{b}}\left[\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{k=1}^{T}c_{k}^{m}\left(a_{k+1}^{m}\right)|\bm{r}_{1},\bm{s}_{1},\bm{a}_{1},\mathcal{H}_{0}\right]\right]$
$=\mathbb{E}\left[\mathbb{E}_{\bm{\pi}}\left[\sum_{m=1}^{M}\sum_{k=1}^{T}c_{k}^{m}\left(a_{k+1}^{m}\right)|\mathcal{H}_{1}\right]-\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{k=1}^{T}c_{k}^{m}\left(a_{k+1}^{m}\right)|\mathcal{H}_{1}\right]\right]$
$+\text{Regret}\left(\pi,\mathcal{H}_{0}\right).$
We can continue this trick to $\mathbb{E}_{\bm{\pi}}\left[\sum_{m=1}^{M}\sum_{k=1}^{T}c_{k}^{m}\left(a_{k+1}^{m}\right)|\mathcal{H}_{1}\right]-\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{k=1}^{T}c_{k}^{m}\left(a_{k+1}^{m}\right)|\mathcal{H}_{1}\right]$, this will give us the desired result. ∎
D.3 Proof of Theorem 4.4
See 4.4 Before we prove the theorem, we first provide a technical lemma showing the concentration property of the empirical mean estimator $\hat{I}_{t}^{m}(s)$, which is a direct result from Howard et al. 2021.
Lemma D.2.
For any $\alpha>0$, $0\leq t\leq T$, we have
$\mathbb{P}\left(\left|\hat{I}_{t}^{m}(s)-I^{m}(s)\right|\geq 1.7\sqrt{\frac{\log\log(2N_{s,m}(t))+0.72\log(10.4/\alpha)}{N_{s,m}(t)}}\right)\leq\alpha.$
Proof.
The proof will proceed by first showing that the cumulative regret can be decomposed into per-timestep regret $\text{Regret}_{t}(\bm{\pi},\mathcal{H}_{t})$. Then we will use the standard technique to bound the number of times suboptimal agents are given action 1.
By Lemma D.1, we only need to bound $\text{Regret}_{t}(\pi,\mathcal{H}_{t})$. We have
$\text{Regret}_{t}(\pi,\mathcal{H}_{t})$ $=\mathbb{E}_{\pi_{t+1}}\left[\sum_{m=1}^{M}c_{t}^{m}(a_{t+1}^{m})\mid\mathcal{H}_{t}\right]+\mathbb{E}_{\pi_{t+1}}\left[\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{k=t+1}^{T}c_{k}^{m}\left(a_{t+1}^{m}\right)|\bm{r}_{t+1},\bm{s}_{t+1},\bm{a}_{t+1},\mathcal{H}_{t}\right]\right]$
$-\mathbb{E}_{\pi_{t+1}^{b}}\left[\sum_{m=1}^{M}c_{t}^{m}\left(a_{t+1}^{m}\right)\mid\mathcal{H}_{t}\right]-\mathbb{E}_{\pi_{t+1}^{b}}\left[\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{k=t+1}^{T}c_{k}^{m}\left(a_{t+1}^{m}\right)|\bm{r}_{t+1},\bm{s}_{t+1},\bm{a}_{t+1},\mathcal{H}_{t}\right]\right]$
$=\mathbb{E}_{\pi_{t+1}}\left[\sum_{m=1}^{M}c_{t}^{m}(a_{t+1}^{m})\mid\mathcal{H}_{t}\right]+\mathbb{E}_{\pi_{t+1}}\left[\mathbb{E}_{\bm{\pi}^{b}}\left[\sum_{m=1}^{M}\sum_{k=t+1}^{T}c_{k}^{m}\left(a_{t+1}^{m}\right)|\bm{r}_{t+1},\bm{s}_{t+1},\bm{a}_{t+1},\mathcal{H}_{t}\right]\right]$
$\leq\mathbb{E}\left[\sum_{\Delta_{s}^{m}>0}\Delta_{s}^{m}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=1\right\}\right]+\mathbb{E}_{\pi_{t+1}}\left[V_{t+1}^{\bm{\pi}^{b}}\left(\bm{s}_{t+1}\right)\mid\bm{s}_{t+1},\bm{a}_{t+1},\mathcal{H}_{t}\right]$
$=\mathbb{E}\left[\sum_{\Delta_{s}^{m}>0}\Delta_{s}^{m}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=1\right\}\right],$
where the second and the third equality holds by Lemma 4.3. Then we have the cumulative regret is upper bounded by
$\sum_{\Delta_{s}^{m}>0}\Delta_{s}^{m}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s\right\}\right].$
We have
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s\right\}\right]$ $=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\geq\gamma\right\}\right]$ (D.71)
$+\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma\right\}\right].$ (D.72)
We will bound the two terms separately. For Eq. (D.71), we have
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\geq\gamma\right\}\right]$
$=\mathbb{E}\left[\sum_{k=1}^{\infty}\mathbbm{1}\left\{\mathrm{LCB}_{\tau_{k}^{m}(s)-1}^{m}(s)\geq\gamma,\tau_{k}^{m}(s)\leq T\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=1}^{T}\mathbbm{1}\left\{\hat{I}_{\tau_{k}^{m}(s)-1}^{m}(s)\geq\gamma+\sqrt{\frac{3\log\left(N_{s,m}^{\left(\tau_{k}^{m}(s)\right)}\right)}{N_{s,m}^{\left(\tau_{k}^{m}(s)\right)}}}\right\}\right]$
$\leq\sum_{k=1}^{T}\exp\left(-N_{s,m}^{\left(\tau_{k}^{m}(s)\right)}\left(\Delta_{s}^{m}\right)^{2}+\log\log\left(N_{s,m}^{\left(\tau_{k}^{m}(s)\right)}\right)-3\log\left(N_{s,m}^{\left(\tau_{k}^{m}(s)\right)}\right)\right)$
$\leq\sum_{k=1}^{T}\frac{1}{k^{2}}\exp\left(-k\left(\Delta_{s}^{m}\right)^{2}\right)$
$\leq 2,$ (D.73)
where $\tau_{k}^{m}(s):=\inf\{t:N_{s,m}^{(t)}=k\}$ and the second inequality holds because of Lemma D.2.
For Eq. (D.72), we have
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma\right\}\right]$
$=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma,\exists(j,s_{t}^{j})\in\mathcal{G},\mathrm{LCB}_{t}^{j}(s_{t}^{j})\leq\gamma\right\}\right]$
$\leq\mathbb{E}\left[\sum_{t=0}^{T}\sum_{(j,s_{j})\in\mathcal{G}}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma,\mathrm{LCB}_{t}^{j}(s_{j})\leq\gamma,s_{t}^{j}=s_{j}\right\}\right]$
$=\sum_{(j,s_{j})\in\mathcal{G}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma,\mathrm{LCB}_{t}^{j}(s_{j})\leq\gamma,s_{t}^{j}=s_{j}\right\}\right]$
For each $(j,s_{j})\in\mathcal{G}$, we have
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma,\mathrm{LCB}_{t}^{j}(s_{j})\leq\gamma,s_{t}^{j}=s_{j}\right\}\right]$
$=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{j}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma,\mathrm{LCB}_{t}^{j}(s_{j})\leq\gamma,s_{t}^{j}=s_{j}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{j}=1,\mathrm{LCB}_{t}^{j}(s_{j})\leq\gamma,s_{t}^{j}=s_{j}\right\}\right]$
$=\mathbb{E}\left[\sum_{k=1}^{\infty}\mathbbm{1}\left\{\mathrm{LCB}_{\tau_{k}^{j}(s_{j})-1}^{m}(s_{j})\geq\gamma,\tau_{k}^{j}(s_{j})\leq T\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=1}^{T}\mathbbm{1}\left\{\hat{I}_{\tau_{k}^{m}(s_{j})-1}^{m}(s_{j})\geq\gamma+\sqrt{\frac{3\log\left(N_{s_{j},m}^{\left(\tau_{k}^{m}(s_{j})\right)}\right)}{N_{s_{j},m}^{\left(\tau_{k}^{m}(s_{j})\right)}}}\right\}\right]$
$\leq u+\mathbb{E}\left[\sum_{k=u+1}^{T}\mathbbm{1}\left\{\hat{I}_{\tau_{k}^{m}(s_{j})-1}^{m}(s_{j})\geq\gamma+\sqrt{\frac{3\log\left(N_{s_{j},m}^{\left(\tau_{k}^{m}(s_{j})\right)}\right)}{N_{s_{j},m}^{\left(\tau_{k}^{m}(s_{j})\right)}}}\right\}\right]$
$\leq u+\sum_{k=u+1}^{T}\frac{1}{k^{2}}$
$\leq u+2,$
where $u=\frac{12}{\left(\Delta_{s_{\min}}^{m_{\min}}\right)^{2}}\left(\log\left(\frac{e}{\left(\Delta_{s_{\min}}^{m_{\min}}\right)^{6}}\right)+\log\log\left(\frac{1}{\left(\Delta_{s_{\min}}^{m_{\min}}\right)^{6}}\right)\right)$. We should note that such choice of $u$ satisfies $\sqrt{3\log(u)/u}\leq\Delta_{s_{\min}}^{m_{\min}}$ (Lemma 18, Garivier & Kaufmann (2016)). Therefore, Eq. (D.72) can be upper bounded as
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t-1}^{m}(s)\leq\gamma\right\}\right]$
$=\sum_{(j,s_{j})\in\mathcal{G}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t-1}^{m}(s)\leq\gamma,\mathrm{LCB}_{t-1}^{j}(s_{j}),s_{t}^{j}=s_{j}\leq\gamma\right\}\right]$
$\leq|\mathcal{G}|(u+2).$ (D.74)
Combining Eq. (D.73) and Eq. (D.74) will give us the desired result.
Further assume that $k\left|\Delta_{s_{\min}}^{m_{\min}}\right|\geq\Delta_{s}^{m}$, for all $\Delta_{s}^{m}>0$. Then we have
$\mathrm{R}(\pi)$ $\leq\sum_{\Delta_{s}^{m}\leq\Delta}\Delta_{s}^{m}\mathbb{E}\left[T_{s}^{m}\right]+\sum_{\Delta_{s}^{m}\geq\Delta}\frac{24k|\mathcal{G}|}{\Delta_{s_{\min}}^{m_{\min}}}\log\left(\frac{e}{\left(\Delta_{s_{\min}}^{m_{\min}}\right)^{6}}\right)+\sum_{\Delta_{s}^{m}>0}(2|\mathcal{G}|+2)\Delta_{s}^{m}$
$\leq B\Delta T+(|\mathcal{S}|M-|\mathcal{G}|)|\mathcal{G}|(24k^{2})\frac{1}{\Delta}\log\left(\frac{e}{\Delta^{6}}\right)+\sum_{\Delta_{s}^{m}>0}(2|\mathcal{G}|+2)\Delta_{s}^{m}.$
Take $\Delta=\frac{24k^{2}|\mathcal{G}|(|\mathcal{S}|M-|\mathcal{G}|)}{BT}$ will give us the desired result. ∎
Appendix E Implementation Detail
E.1 Regret Curve
We will describe the details of implementation. Specifically, we will randomly generate a number between $[0,1]$ independently and assign this number to be $P_{a}^{m}(s,1),\forall a,m,s\in\{0,1\}$. To make sure Assumption 3.1 hold, we will switch the number if $P_{0}^{m}(s,1)>P_{1}^{m}(s,1)$.
Without loss of generality, assume that $I^{1}(1)\geq I^{1}(0)\geq I^{2}(1)\geq I^{2}(0)$. We choose $\gamma$ to be between $I^{1}(0)$ and $I^{2}(0)$. Specifically, let $\theta\in(0,1)$ be a tunable parameter. We set $\gamma:=\theta I^{2}(1)+(1-\theta)I^{2}(0)$. To make sure that the gap between $\gamma$ and $I^{2}(1)$ is not too small, we choose those instances where $I^{2}(1)-I^{2}(0)\geq 0.15$. To generate the three different $\gamma$ values in Figure 1, we selected three different choices of $\theta$: $0.1,0.5,0.9$, yielding the observed $\Delta_{s_{\min}}^{m_{\min}}$ values.
E.2 Cumulative Reward
Here we will discuss more on the implementation of the algorithms. First of all for the UCWhittle families introduced in Wang et al. (2023b), the algorithms reinitialize the initial state after certain time-steps. We do not include this episodic structure as this contradicts our goal of maximizing the finite-horizon cumulative reward. We keep the rest structures of their algorithms as the same.
We additionally note that the code provided by Wang et al. (2023b) contains errors and could not be used to reproduce the results shown in Wang et al. (2023b). A mistake was found in Line 73 of uc__\__whittle.py that they provided. Specifically, they use the current state of an agent as the index to get the current action of this agent. As a result, we build our simulation environment from scratch.
Jiamin: Furthermore, we note that when the stochatsic dominance assumption does not hold, the Whittle’s index could be negative. Meanwhile, even when the stochastic dominance assumption holds, the estimator of the transition probability given by UCWhittleUCBPriv and UCWhittlePvPriv is not guaranteed to satsify the stochastic dominance assumption. Therefore, we make modifications to the binary search rountine of getting Whittle’s index. Specifically, we let the initial lower bound be $-10$ to ensure a more accurate calculation of Whittle’s index.
Adversarial Setting
Figure E.8: Average cumulative reward under $B=5$ (left column), $10$ (middle column), $20$ (right column). Averaged over $50$ instance satisfying the opposite of Assumption 3.1, each with $13$ repetitions. Noiseless reward.
In this paragraph, we will modify the generated instances to make sure that not giving intervention will stochastically dominate giving intervention. After an instance is generated, we check whether there exists $s,m$ such that $P_{1}^{m}(s,1)>P_{0}^{m}(s,1)$. If there does, we will simply swap these two transitions. We will test the performance of LGGT comparing to the same sets of policies.
In Figure E.8, we notice that LGGT with threshold $\gamma=0$ performs similarly to UCWhittlePvPriv and is better than other online algorithms for all three budgets considered across 1000 horizons. We emphasize that in this setting, the optimal policy would be not giving any intervention at all, which makes the problem easy.